### v1

# A Two-Stage Fourth-Order Implicit Scheme for Stiff problems

Abstract
This paper presents a novel two-stage fourth-order implicit scheme designed to overcome the limitations of existing spatiotemporal coupling two-stage fourth-order methods, which have thus far been confined to explicit frameworks. In such frameworks, computational efficiency is severely constrained by the CFL condition when addressing stiff problems, and they fundamentally conflicts with spatiotemporal coupled methods whose core advantage lies in embedding stiff source terms. By fully leveraging temporal derivatives of physical variables within a rigorously derived compact mathematical framework, the scheme achieves fourth-order temporal accuracy in only two stages. Furthermore, a sufficient condition for A-stability is established through systematic theoretical and numerical investigation, and a Newton iterative procedure is provided to accelerate convergence. Extensive numerical experiments on classical stiff problems confirm the method’s effectiveness and competitiveness.
Key words: two-stage fourth-order; temporal derivatives; implicit scheme; A-stability; stiff problems; Newton iteration
1 Introduction
The generalized Riemann problem (GRP) fundamentally differs from the classical Riemann problem through its use of spatially varying initial data, which generates non-self-similar solutions with non-zero time derivatives. This critical advancement enables the natural incorporation of physical source terms and other essential effects (e.g., continuous entropy variation) into the temporal evolution, establishing GRP as an indispensable framework for capturing complex physical mechanisms. The simplest GRP form replaces piecewise constant with piecewise linear initial data in homogeneous systems, first explored by Kolgan [1] and further developed by van Leer [2, 3]. Toro and Titarev [4] later extended GRP to include source terms using the Lax-Wendroff procedure, forming the basis of arbitrarily high-order (ADER) schemes [5, 6] applicable to hydrodynamics, magnetohydrodynamics [7], reaction-diffusion [8], and stiff systems [9], though these linearized solvers struggle under extreme thermodynamic conditions. Ben-Artzi et al. [10, 11, 12] were the first to develop the GRP solver for compressible fluids, which can be considered as the nonlinear version of the ADER solver. By incorporating thermodynamic effects, the GRP solver can accurately resolve strong waves, making it valuable under extreme conditions [13]. The GRP solver has since been successfully applied to duct flow [14], reactive flow [15], shallow water equations [16, 17], multi-material and multiphase flow [18, 19, 20, 21], magnetohydrodynamics [22], relativistic hydrodynamics [23, 24, 25], blood flow modeling [26], and radiation hydrodynamics [27].
The inherent spatiotemporal coupling in fluid dynamics necessitates numerical algorithms that preserve such correlations, which conventional Riemann solvers fail to achieve as they only provide piecewise constant states for constructing instantaneous Cauchy fluxes. This limitation leads to error accumulation over time in flows with strong discontinuities. In contrast, generalized Riemann solvers (e.g., ADER or GRP) provide temporal derivatives of physical quantities, enabling construction of time-interval fluxes with inherent spatiotemporal coupling. These fluxes exhibit Lipschitz continuity with respect to boundary perturbations, allowing temporal effects to enhance spatial regularity. Consequently, errors from strong discontinuities can be directly controlled through the time step, enabling accurate long-term resolution while establishing consistency between mathematical weak solutions and physical integral balance laws [28, 29].
Runge-Kutta methods, originally developed by Carl Runge [30] in 1895 and Martin Wilhelm Kutta [31] in 1901, constitute a family of classical numerical techniques for solving initial-value problems of ordinary differential equations that comprise both explicit and implicit forms. The explicit variants, exemplified by the classical fourth-order RK4 scheme, achieve high-order accuracy through weighted averages of function values at multiple stages, offering implementation simplicity without requiring calculation of higher-order derivatives. However, their stability remains constrained by the CFL condition, rendering them inefficient for stiff problems. In contrast, Implicit Runge-Kutta methods proposed by Kuntzmann [32] and Butcher [33], such as Radau and Gauss-type schemes, compute slopes at future time steps by solving systems of nonlinear equations. Although computationally more expensive, they exhibit superior numerical stability, making them particularly suitable for stiff differential equations.
Recently, Li et al. developed a two-stage fourth-order time advancement method [34], which makes full use of the temporal derivatives of physical quantities, enabling the scheme to achieve fourth-order time accuracy in only two stages. This method successfully combines the simplicity advantages of Runge-Kutta-type methods with the spatiotemporally coupling properties of the GRP solver. Then they details the characteristics of this class of methods [35]: (i) Computational efficiency: Under the same conditions, its computational efficiency is twice that of the Runge-Kutta method, which requires four stages to achieve fourth-order time accuracy [36]; (ii) Compactness: Since the number of time advancement steps is reduced by half, the stencil size is also halved; the spatial-temporal coupled HWENO-type reconstruction [37, 38] can make the computational scheme more compact; (iii) Stability: It has been proven that its stability region is larger than that of Runge-Kutta methods [39]; (iv) Compatibility: The method can be easily combined with other methods, such as the GKS solver [40], the multi-moment method [41], and the CRP reconstruction technique [42].
However, despite the aforementioned advantages, current spatiotemporally coupled two-step fourth-order methods still suffer from significant limitations. On the one hand, such methods are currently confined to explicit frameworks. As a result, their time step is severely constrained by the CFL stability condition, and this restriction in turn drastically reduces their computational efficiency when dealing with stiff problems. On the another hand, the stiffness of a problem is often predominantly manifested in its source terms. For instance, in problems involving multi-scale physics like combustion or chemically reacting flows, the source terms can introduce significant stiffness. The key advantage of the GRP solver over the conventional Riemann solver lies in its ability to provide the temporal derivatives of physical quantities, and this derivative information inherently encodes the effects of the source terms from the governing equations. However, a critical paradox emerges: this same temporal derivative data is utilized in two distinct ways: it is essential for constructing the time-interval fluxes and serves as the cornerstone for building the two-stage fourth-order explicit scheme. The former (time-interval fluxes) can successfully incorporate information from stiff source terms, whereas the latter (the explicit scheme itself) is fundamentally unsuitable for solving stiff problems due to its explicit nature. This inherent incompatibility points to a significant limitation in the current methodological framework. Inaddition, the incorporation of temporal derivatives may allows the proposed method to achieve not only higher computational accuracy but also a more compact formulation compared to existing implicit schemes. Based on the above considerations, we conclude that there is an urgent need to develop two-stage fourth-order implicit methods to address multi-scale problems such as those involving strong stiffness.
As the foundational component of this research on spatiotemporally coupled two-stage fourth-order implicit schemes, hereafter referred to as Implicit TSFO, we develop an implicit framework that attains fourth-order accuracy in two stages through the incorporation of temporal derivatives of physical quantities. Section 2 presents the detailed formulation of the scheme via rigorous mathematical derivation. This is followed in Section 3 by a thorough analysis of the stability conditions of the implicit method. Section 4 outlines the implementation procedure based on Newton iteration. In Section 5, the efficacy of the proposed scheme is demonstrated through several canonical numerical examples, supplemented by comparative studies with classical fourth-order implicit and explicit Runge-Kutta methods. Finally, Section 6 concludes the paper and suggests potential directions for future research.
2 The Construction of the Implicit TSFO Scheme
Consider the time-dependent equations
$\frac{\partial\bf u}{\partial t}=\mathcal{L}(\bf u),$ (2.1)
where we assume that the operator $\mathcal{L}$ is sufficiently smooth, meaning all required derivatives exist. We write (2.1) as
${\bf u}^{n+\theta}={\bf u}^{n}+\int_{t_{n}}^{t_{n}+\theta\Delta t}\mathcal{L}[{\bf u}(t)]dt,$ (2.2)
where ${\bf u}^{n+\theta}={\bf u}(t_{n}+\theta\Delta t)$, $\theta\in[0,1]$, denote $\mathcal{Q}(t)=\mathcal{L}[{\bf u}(t)]$, then
$\mathcal{Q}^{\prime}(t)=\mathcal{L}_{\bf u}\mathcal{L},\quad\mathcal{Q}^{\prime\prime}(t)=\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L},\quad\mathcal{Q}^{\prime\prime\prime}(t)=\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L},$ (2.3)
and
$\int_{t_{n}}^{t_{n}+\theta\Delta t}\mathcal{L}[{\bf u}(t)]dt$ (2.4)
$=\int_{t_{n}}^{t_{n}+\theta\Delta t}\mathcal{Q}(t)dt$
$=\int_{t_{n}}^{t_{n}+\theta\Delta t}\left\{\mathcal{Q}(t_{n})+\frac{(t-t_{n})}{1!}\mathcal{Q}^{\prime}(t_{n})+\frac{(t-t_{n})^{2}}{2!}\mathcal{Q}^{\prime\prime}(t_{n})+\frac{(t-t_{n})^{3}}{3!}\mathcal{Q}^{\prime\prime\prime}(t_{n})+\mathcal{O}[(t-t_{n})^{4}]\right\}dt$
$=\frac{\theta\Delta t}{1!}\mathcal{Q}(t_{n})+\frac{(\theta\Delta t)^{2}}{2!}\mathcal{Q}^{\prime}(t_{n})+\frac{(\theta\Delta t)^{3}}{3!}\mathcal{Q}^{\prime\prime}(t_{n})+\frac{(\theta\Delta t)^{4}}{4!}\mathcal{Q}^{\prime\prime\prime}(t_{n})+\mathcal{O}[(\theta\Delta t)^{5}].$
Especially for $\theta=\frac{1}{2}$ and 1, we have
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{\Delta t}{2}\mathcal{L}({\bf u}^{n})+\frac{(\Delta t)^{2}}{8}[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})+\frac{(\Delta t)^{3}}{48}[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$ (2.5)
$\quad\quad\quad\quad\quad\;\;+\frac{(\Delta t)^{4}}{384}[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})+\mathcal{O}(\Delta t^{5}),$
and
${\bf u}^{n+1}={\bf u}^{n}+\Delta t\mathcal{L}({\bf u}^{n})+\frac{(\Delta t)^{2}}{2}[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})+\frac{(\Delta t)^{3}}{6}[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$ (2.6)
$\quad\quad\quad\quad\;\;\;+\frac{(\Delta t)^{4}}{24}[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})+\mathcal{O}(\Delta t^{5}).$
We construct a two-stage fourth-order implicit scheme by utilizing the values of physical quantities along with their temporal derivatives. The proposed scheme is formulated as follows:
Stage 1. We construct the following implicit scheme by utilizing the physical quantities at the time level $t_{n}$ and $t_{n+\frac{1}{2}}$ to obtain the solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$,
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\Delta t\left[A_{1}\mathcal{L}({\bf u}^{n})+A_{2}\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]+\Delta t^{2}\left[B_{1}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+B_{2}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right],$ (2.7)
where the coefficients $A_{1}$, $A_{2}$, $B_{1}$, $B_{2}$ are to be determined to achieve fourth-order accuracy in Stage 1.
Stage 2. We construct the following implicit scheme by utilizing the physical quantities at the time level $t_{n}$, $t_{n+\frac{1}{2}}$ and $t_{n+1}$ to obtain the solution at the next time level ${\bf u}^{n+1}$,
${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left[A_{3}\mathcal{L}({\bf u}^{n})+A_{4}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+A_{5}\mathcal{L}({\bf u}^{n+1})\right]$ (2.8)
$+\Delta t^{2}\left[B_{3}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+B_{4}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+B_{5}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right],$
where the coefficients $A_{3}$, $A_{4}$, $A_{5}$, $B_{3}$, $B_{4}$, $B_{5}$ are to be determined to achieve fourth-order accuracy in $\bf Stage2$.
By treating ${\bf u}^{n}$, ${\bf u}^{n+\frac{1}{2}}$, and ${\bf u}^{n+1}$ as the exact solution and using Taylor series expansions, we obtain
${\bf u}^{n+\theta}={\bf u}^{n}+\frac{\theta\Delta t}{1!}\frac{\partial{\bf u}}{\partial t}{\Big\lvert}_{t_{n}}+\frac{(\theta\Delta t)^{2}}{2!}\frac{\partial^{2}{\bf u}}{\partial t^{2}}{\Big\lvert}_{t_{n}}+\frac{(\theta\Delta t)^{3}}{3!}\frac{\partial^{3}{\bf u}}{\partial t^{3}}{\Big\lvert}_{t_{n}}+\mathcal{O}[(\theta\Delta t)^{4}],$ (2.9)
where
$\frac{\partial\bf u}{\partial t}=\mathcal{L},\quad\frac{\partial^{2}\bf u}{\partial t^{2}}=\mathcal{L}_{\bf u}\mathcal{L},\quad\frac{\partial^{3}\bf u}{\partial t^{3}}=\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}.$ (2.10)
Especially for $\theta=\frac{1}{2}$ and 1, we have
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{\Delta t}{2}\mathcal{L}({\bf u}^{n})+\frac{(\Delta t)^{2}}{8}[\mathcal{L}_{u}\mathcal{L}]({\bf u}^{n})+\frac{(\Delta t)^{3}}{48}[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})+\mathcal{O}[(\Delta t)^{4}],$ (2.11)
and
${\bf u}^{n+1}={\bf u}^{n}+\Delta t\mathcal{L}({\bf u}^{n})+\frac{(\Delta t)^{2}}{2}[\mathcal{L}_{u}\mathcal{L}]({\bf u}^{n})\\ +\frac{(\Delta t)^{3}}{6}[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})+\mathcal{O}[(\Delta t)^{4}].$ (2.12)
Taylor-expand the operators $\mathcal{L}$ with respect to $\bf u$, yielding
$\mathcal{L}({\bf u}^{n+\theta})$ $=\mathcal{L}[{\bf u}^{n}+({\bf u}^{n+\theta}-{\bf u}^{n})]$ (2.13)
$=\mathcal{L}({\bf u}^{n})+\frac{{\bf u}^{n+\theta}-{\bf u}^{n}}{1!}\mathcal{L}_{\bf u}({\bf u}^{n})+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2!}\mathcal{L}_{\bf uu}({\bf u}^{n})$
$\quad\quad\quad\quad+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{3}}{3!}\mathcal{L}_{\bf uuu}({\bf u}^{n})+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{4}]$
$=\mathcal{L}({\bf u}^{n})+({\bf u}^{n+\theta}-{\bf u}^{n})\mathcal{L}_{\bf u}({\bf u}^{n})+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2}\mathcal{L}_{\bf uu}({\bf u}^{n})$
$\quad\quad\quad\quad+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{3}}{6}\mathcal{L}_{\bf uuu}({\bf u}^{n})+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{4}].$
Denote $\mathcal{G}=\frac{\partial}{\partial t}\mathcal{L}$, then
$\mathcal{G}=\mathcal{L}_{\bf u}\mathcal{L},\quad\mathcal{G}_{\bf u}=\mathcal{L}_{\bf uu}\mathcal{L}+\mathcal{L}_{\bf u}^{2},\quad\mathcal{G}_{\bf uu}=\mathcal{L}_{\bf uuu}\mathcal{L}+3\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}.$ (2.14)
Taylor-expand the operators $\mathcal{G}$ with respect to $\bf u$, yielding
$\mathcal{G}({\bf u}^{n+\theta})$ $=\mathcal{G}[{\bf u}^{n}+({\bf u}^{n+\theta}-{\bf u}^{n})]$ (2.15)
$=\mathcal{G}({\bf u}^{n})+\frac{{\bf u}^{n+\theta}-{\bf u}^{n}}{1!}\mathcal{G}_{\bf u}({\bf u}^{n})+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2!}\mathcal{G}_{\bf uu}({\bf u}^{n})+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{3}]$
$=[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})+({\bf u}^{n+\theta}-{\bf u}^{n})[\mathcal{L}_{\bf uu}\mathcal{L}+\mathcal{L}_{\bf u}^{2}]({\bf u}^{n})$
$\quad\quad\quad\quad\quad\quad+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2}[\mathcal{L}_{\bf uuu}\mathcal{L}+3\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}]+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{3}].$
Substituting (2.11), (2.13), (2.15) into (2.7), we obtain
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\Delta t\left(A_{1}+A_{2}\right)\mathcal{L}({\bf u}^{n})$ (2.16)
$\quad\quad\quad\quad\quad\;\;+(\Delta t)^{2}\left(\frac{1}{2}A_{2}+B_{1}+B_{2}\right)[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\quad\;\;+(\Delta t)^{3}\left(\frac{1}{8}A_{2}+\frac{1}{2}B_{2}\right)[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\quad\;\;+(\Delta t)^{4}\left(\frac{1}{48}A_{2}+\frac{1}{8}B_{2}\right)[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\quad\;\;+\mathcal{O}[(\Delta t)^{5}].$
Compare (2.5) with (2.16), we obtain that the coefficients must satisfy the following equation relationship,
$\begin{cases}A_{1}+A_{2}=\frac{1}{2},\\ \frac{1}{2}A_{2}+B_{1}+B_{2}=\frac{1}{8},\\ \frac{1}{8}A_{2}+\frac{1}{2}B_{2}=\frac{1}{48},\\ \frac{1}{48}A_{2}+\frac{1}{8}B_{2}=\frac{1}{384},\end{cases}$ (2.17)
Solve the above system of equations, we obtain the unique solution
$A_{1}=\frac{1}{4},\quad A_{2}=\frac{1}{4},\quad B_{1}=\frac{1}{48},\quad B_{2}=-\frac{1}{48}.$ (2.18)
Substituting (2.11), (2.12), (2.13) and (2.15) into (2.8), we obtain
${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left(A_{3}+A_{4}+A_{5}\right)\mathcal{L}({\bf u}^{n})$ (2.19)
$+\Delta t^{2}\left(\frac{1}{2}A_{4}+A_{5}+B_{3}+B_{4}+B_{5}\right)[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})$
$+\Delta t^{3}\left(\frac{1}{8}A_{4}+\frac{1}{2}A_{5}+\frac{1}{2}B_{4}+B_{5}\right)[\mathcal{L}_{\bf u}^{2}\mathcal{L}+\mathcal{L}_{\bf uu}\mathcal{L}^{2}]({\bf u}^{n})$
$+\Delta t^{4}\left(\frac{1}{48}A_{4}+\frac{1}{6}A_{5}+\frac{1}{8}B_{4}+\frac{1}{2}B_{5}\right)[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})$
$+\mathcal{O}(\Delta t^{5})$
The comparison of (2.6) and (2.19), gives
$\begin{cases}A_{3}+A_{4}+A_{5}=1,\\ \frac{1}{2}A_{4}+A_{5}+B_{3}+B_{4}+B_{5}=\frac{1}{2},\\ \frac{1}{8}A_{4}+\frac{1}{2}A_{5}+\frac{1}{2}B_{4}+B_{5}=\frac{1}{6},\\ \frac{1}{48}A_{4}+\frac{1}{6}A_{5}+\frac{1}{8}B_{4}+\frac{1}{2}B_{5}=\frac{1}{24},\end{cases}$ (2.20)
Solve the above system of equations, we obtain
$\begin{cases}A_{3}=\frac{1}{6}+4C+\frac{1}{2}D,\\ A_{4}=\frac{2}{3}-8C+2D,\\ A_{5}=\frac{1}{6}+4C-\frac{5}{2}D,\\ B_{3}=C,\\ B_{4}=D,\\ B_{5}=\frac{1}{2}D-C.\end{cases}$ (2.21)
where $C$, $D$ are real parameters to be determined. Substituting (2.18) and (2.21) into (2.7) and (2.8) respectively, we obtain the specific form of the desired two-stage fourth-order implicit (e.g., Implicit TSFO) scheme as follows:
Stage 1. Iteratively solve the following implicit scheme to obtain the solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$,
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{\Delta t}{4}\left[\mathcal{L}({\bf u}^{n})+\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]-\frac{\Delta t^{2}}{48}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right].$ (2.22)
Stage 2. Iteratively solve the following implicit scheme to obtain the solution at the next time level ${\bf u}^{n}$,
${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left[\left(\frac{1}{6}+4C+\frac{1}{2}D\right)\mathcal{L}({\bf u}^{n})+\left(\frac{2}{3}-8C+2D\right)\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{6}+4C-\frac{5}{2}D\right)\mathcal{L}({\bf u}^{n+1})\right]$ (2.23)
$+\Delta t^{2}\left[C\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+D\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{2}D-C\right)\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right].$
3 The Stability Analysis of the Implicit TSFO Scheme
To examine the stability of the implicit scheme (2.22)-(2.23), we consider the following model equation
$\mathcal{L}(\bf u)=\lambda\bf u,$ (3.1)
where $\lambda$ is the eigenvalue of the Jacobian matrix of the system (2.1), and ${\rm Re}(\lambda)\leq 0$. Combining (3.1) and (2.1), we obtain
$\frac{\partial}{\partial t}\mathcal{L}({\bf u})=\lambda^{2}{\bf u}.$ (3.2)
Subtitute (3.1)-(3.2) into the implicit scheme (2.22)-(2.23), we obtain
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{z}{4}\left({\bf u}^{n}+{\bf u}^{n+\frac{1}{2}}\right)-\frac{z^{2}}{48}\left({\bf u}^{n+\frac{1}{2}}-{\bf u}^{n}\right),$ (3.3)
${\bf u}^{n+1}={\bf u}^{n}$ $+z\left[\left(\frac{1}{6}+4C+\frac{1}{2}D\right){\bf u}^{n}+\left(\frac{2}{3}-8C+2D\right){\bf u}^{n+\frac{1}{2}}+\left(\frac{1}{6}+4C-\frac{5}{2}D\right){\bf u}^{n+1}\right]$ (3.4)
$+z^{2}\left[C{\bf u}^{n}+D{\bf u}^{n+\frac{1}{2}}+\left(\frac{1}{2}D-C\right){\bf u}^{n+1}\right],$
where $z=\lambda\Delta t$. From (3.3), we obtain
${\bf u}^{n+\frac{1}{2}}=R(z){\bf u}^{n},$ (3.5)
where
$R(z)=\frac{1+\frac{z}{4}+\frac{z^{2}}{48}}{1-\frac{z}{4}+\frac{z^{2}}{48}}=1+\frac{24}{z}+\mathcal{O}\left(\frac{1}{z^{2}}\right).$ (3.6)
Substitute (3.5) into (3.4), we obtain
${\bf u}^{n+1}=G(z){\bf u}^{n},$ (3.7)
where
$G(z)$ $=\frac{1+z\left[\left(\frac{1}{6}+4C+\frac{1}{2}D\right)+\left(\frac{2}{3}-8C+2D\right)R(z)\right]+z^{2}\left[C+DR(z)\right]}{1-\left(\frac{1}{6}+4C-\frac{5}{2}D\right)z-\left(\frac{1}{2}D-C\right)z^{2}}$ (3.8)
$\sim\frac{C+D}{C-\frac{1}{2}D},\quad z\to\infty.$
Therefore, we have the following crucial conclusion.
Lemma 3.1.
The amplification factor $G$ satisfying
$G(\infty)=0,$ (3.9)
if and only if
$C+D=0,$ (3.10)
where $C$ and $D$ are the undetermined parameters in the Implicit TSFO scheme (2.22)-(2.23).
Under the condition (3.10), the amplification factor (3.8) reduces to
$G(z)=\frac{1+z\left[\frac{1}{6}+\frac{7}{2}C+\left(\frac{2}{3}-10C\right)R(z)\right]+Cz^{2}\left[1-R(z)\right]}{1-\left(\frac{1}{6}+\frac{13}{2}C\right)z+\frac{3}{2}Cz^{2}}.$ (3.11)
Our objective in the following is to determine the range of $C$ for which the Implicit TSFO scheme (2.22)-(2.23) is A-stable. Achieving this requires not only need condition (3.9) above but also need the following condition:
$\sup_{\text{Re}(z)\leq 0}|G(z)|\leq 1.$ (3.12)
The theoretical foundation for verifying A-stability is based on the maximum modulus principle: If $G(z)$ is analytic in the closed left half-plane $\text{Re}(z)\leq 0$, then the maximum of $\lvert G(z)\lvert$ occurs on the boundary, specifically on: the imaginary axis $z=iy,\;y\in\mathbb{R}$ and at infinity $z\to\infty$. Therefore, to achieve the A-stability of the Implicit TSFO scheme (2.22)-(2.23), we need to examine the following three conditions:
Condition 1: $G(z)$ is analytic in the closed left half-plane, i.e.,
$G\in\mathcal{H}(\mathbb{C}_{-}),$ (3.13)
where $\mathbb{C}_{-}=\{z\in\mathbb{C}:\mathrm{Re}(z)\leq 0\}$.
Condition 2: The $L_{\infty}$-norm of $G$ on the imaginary axis is bounded by 1, i.e.,
$\|G(i\cdot)\|_{L_{\infty}(\mathbb{R})}\leq 1,$ (3.14)
or equivalently,
$\sup_{y\in\mathbb{R}}|G(iy)|\leq 1.$ (3.15)
Condition 3: $G$ is bounded at infinity with magnitude not exceeding 1, i.e.,
$|G(\infty)|\leq 1.$ (3.16)
More precisely, the limit
$\lim_{|z|\to\infty,\,\mathrm{Re}(z)\leq 0}G(z)$ (3.17)
exists and its modulus is bounded by 1.
Since Condition 3 (3.16)-(3.17) is already satisfied by Lemma 3.1, we focus on analyzing Conditions 1 and Condition 2 to determine the range of $C$.
Examination of Condition 1, i.e., (3.13).
The first requirement for A-stability is that the amplification factor $G(z)$ must be analytic in the closed left half-plane $\mathbb{C}_{-}$. This is equivalent to requiring that all poles of $G(z)$ lie strictly in the right half-plane. The denominator of $G(z)$ is given by:
$D(z)=\left[1-\left(\frac{1}{6}+\frac{13}{2}C\right)z+\frac{3}{2}Cz^{2}\right]\left(1-\frac{z}{4}+\frac{z^{2}}{48}\right).$ (3.18)
The poles are the roots of $D(z)=0$. The equation
$1-\frac{z}{4}+\frac{z^{2}}{48}=0,$ (3.19)
has two complex roots
$z=6+2i\sqrt{3},\quad z=6-2i\sqrt{3},$ (3.20)
both located in the right half-plane. The roots of the equation
$1-\left(\frac{1}{6}+\frac{13}{2}C\right)z+\frac{3}{2}Cz^{2}=0$ (3.21)
are given by:
$z=\begin{cases}\frac{\frac{1}{3}+13C\pm\sqrt{\left(\frac{1}{3}+13C\right)^{2}-24C}}{6C},&C\neq 0,\\ 6,&C=0.\end{cases}$ (3.22)
So that the real part of $z$ is
$\text{Re}(z)=\begin{cases}\frac{\frac{1}{3}+13C}{6C},&C\neq 0,\\ 6,&C=0.\end{cases}$ (3.23)
Therefore, we have the following conclusion.
Lemma 3.2.
If and only if
$C\geq 0\quad\text{or}\quad C<-\frac{1}{39},$ (3.24)
we have
${\textnormal{Re}}(z)>0,$ (3.25)
i.e., all roots of the equation (3.21) lie in the right half-plane, in another words, the amplification factor $G(z)$ is analytic in $\textnormal{Re}(z)\leq 0$.
Examination of Condition 2, i.e., (3.14)-(3.15).
Substitute $z=iy$ into $R(z)$, we obtain
$R(iy)=\frac{1+i\frac{y}{4}-\frac{y^{2}}{48}}{1-i\frac{y}{4}-\frac{y^{2}}{48}}.$ (3.26)
Let
$N_{R}=1+i\frac{y}{4}-\frac{y^{2}}{48},\quad D_{R}=1-i\frac{y}{4}-\frac{y^{2}}{48}.$ (3.27)
Since $D_{R}=\overline{N_{R}}$ (complex conjugate), we have:
$\lvert R(iy)\lvert=\Big\lvert\frac{N_{R}}{D_{R}}\Big\lvert=\Big\lvert\frac{N_{R}}{\overline{N_{R}}}\Big\lvert=1.$ (3.28)
Thus
$R(iy)=e^{i\psi(y)},$ (3.29)
with phase
$\psi(y)=2\text{arg}\left(1+i\frac{y}{4}-\frac{y^{2}}{48}\right)=2\text{arctan}\left(\frac{y/4}{1-y^{2}/48}\right).$ (3.30)
Substitute $z=iy$ and (3.29) into $G(z)$, we obtain
$G(iy)=\frac{1+iy[\frac{1}{6}+\frac{7}{2}C+(\frac{2}{3}-10C)e^{i\psi}]-Cy^{2}(1-e^{i\psi})}{1-(\frac{1}{6}+\frac{13}{2}C)iy-\frac{3}{2}Cy^{2}}.$ (3.31)
Analytically proving $\lvert G(iy)\lvert\leq 1$ for all $y\in\mathbb{R}$ is complex. Instead, we employ a systematic numerical scanning algorithm to determine the A-stability conditions of the scheme. The parameter space $C\in[0.00,0.10]$ is discretized into 5,000 sampling points, while the frequency domain along the imaginary axis $y\in[10^{-8},10^{4}]$ is densely sampled using 25,000 logarithmically spaced points. For each parameter-frequency pair $(C,y)$, the corresponding amplification factor $\lvert G(iy)\lvert$ is computed. The algorithm identifies the maximum amplification factor across all frequencies and verifies its compliance with the stability criterion $\lvert G(iy)\lvert\leq 1$. The corresponding numerical results are shown in the figure 3.1. Based on the above systematic numerical scanning analysis, Lemma 3.1. and Lemma 3.2., we draw the following conclusion regarding the A-stability of the scheme.
Conclusion 3.3.
The Implicit TSFO scheme, given by equations (2.22)-(2.23), achieves A-stability under the condition
D=C,D=-C, (3.32)
with the parameter $C$ confined to the interval
$C\in[0.018824,0.045589],$ (3.33)
where the optimal parameter
$C=0.018924$ (3.34)
provides superior numerical damping characteristics.
Figure 3.1: A-stability: Valid C Range Determination
By applying the Conclusion 3.3. to the Implicit TSFO scheme (2.22)-(2.23), we derive the final form of this scheme, which is A-stable.
Stage 1. Iteratively solve the following implicit scheme to obtain the solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$,
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{\Delta t}{4}\left[\mathcal{L}({\bf u}^{n})+\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]-\frac{\Delta t^{2}}{48}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right].$ (3.35)
Stage 2. Iteratively solve the following implicit scheme to obtain the solution at the next time level ${\bf u}^{n}$,
${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left[\left(\frac{1}{6}+\frac{7}{2}C\right)\mathcal{L}({\bf u}^{n})+\left(\frac{2}{3}-10C\right)\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{6}+\frac{13}{2}C\right)\mathcal{L}({\bf u}^{n+1})\right]$ (3.36)
$+C\Delta t^{2}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{3}{2}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right],$
where $C\in[0.018824,0.045589]$, with $C=0.018924$ being the optimal value.
4 The Implementation of the Implicit TSFO Scheme
To achieve accelerated convergence, the implicit scheme can be realized through Newton’s iterative method, where the initial guess for each stage uses the result of the two-stage fourth-order explicit scheme [34].
Stage 1. The solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$ is computed iteratively as follows:
  1. 1.
    Initialization:
    $({\bf u}^{n+\frac{1}{2}})^{(0)}={\bf u}^{n}+\frac{\Delta t}{2}\mathcal{L}({\bf u}^{n})+\frac{\Delta t^{2}}{8}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})$ (4.1)
  2. 2.
    Iteration: For $k=0,1,2,\dots$, repeat the following steps until convergence:
    1. (a)
      Solve the linear system for $\Delta{\bf u}$:
      $J_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]\Delta{\bf u}=-F_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]$ (4.2)
    2. (b)
      Update the solution:
      $({\bf u}^{n+\frac{1}{2}})^{(k+1)}=({\bf u}^{n+\frac{1}{2}})^{(k)}+\Delta{\bf u}$ (4.3)
  3. 3.
    Convergence check: The iteration continues while
    $\left\lVert({\bf u}^{n+\frac{1}{2}})^{(k+1)}-({\bf u}^{n+\frac{1}{2}})^{(k)}\right\rVert\geq\mathrm{TOL}$ (4.4)
  4. 4.
    Final assignment: Once converged, set
    ${\bf u}^{n+\frac{1}{2}}=({\bf u}^{n+\frac{1}{2}})^{(k+1)}$ (4.5)
where the Jacobian $J_{1}$ and residual $F_{1}$ are defined as:
$J_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]$ $=\frac{\Delta t}{4}\mathcal{L}_{\bf u}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-\frac{\Delta t^{2}}{48}\left(\frac{\partial}{\partial t}\mathcal{L}\right)_{\bf u}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-I,$ (4.6)
$F_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]$ $={\bf u}^{n}+\frac{\Delta t}{4}\left\{\mathcal{L}{({\bf u}^{n})}+\mathcal{L}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]\right\}$
$\quad-{\frac{\Delta t^{2}}{48}}\left\{\frac{\partial}{\partial t}\mathcal{L}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right\}-({\bf u}^{n+\frac{1}{2}})^{(k)}.$
Stage 2. The solution at the next time level ${\bf u}^{n+1}$ is computed iteratively as follows:
  1. 1.
    Initialization:
    $({\bf u}^{n+1})^{(0)}={\bf u}^{n}+\Delta t\mathcal{L}({\bf u}^{n})+\frac{\Delta t^{2}}{6}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+2\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]$ (4.7)
  2. 2.
    Iteration: For $k=0,1,2,\dots$, repeat the following steps until convergence:
    1. (a)
      Solve the linear system for $\Delta{\bf u}$:
      $J_{2}\left[({\bf u}^{n+1})^{(k)}\right]\Delta{\bf u}=-F_{2}\left[({\bf u}^{n+1})^{(k)}\right]$ (4.8)
    2. (b)
      Update the solution:
      $({\bf u}^{n+1})^{(k+1)}=({\bf u}^{n+1})^{(k)}+\Delta{\bf u}$ (4.9)
  3. 3.
    Convergence check: The iteration continues while
    $\left\lVert\lvert({\bf u}^{n+1})^{(k+1)}-({\bf u}^{n+1})^{(k)}\right\rVert\geq\mathrm{TOL}$ (4.10)
  4. 4.
    Final assignment: Once converged, set
    ${\bf u}^{n+1}=({\bf u}^{n+1})^{(k+1)}$ (4.11)
where the Jacobian $J_{2}$ and residual $F_{2}$ are defined as:
$J_{2}\left[({\bf u}^{n+1})^{(k)}\right]=\Delta t\left(\frac{1}{6}+\frac{13}{2}C\right){\mathcal{L}}_{\bf u}\left[({\bf u}^{n+1})^{(k)}\right]-\frac{3}{2}C\Delta t^{2}\left(\frac{\partial}{\partial t}\mathcal{L}\right)_{\bf u}\left[({\bf u}^{n+1})^{(k)}\right]-I,$ (4.12)
$F_{2}\left[({\bf u}^{n+1})^{(k)}\right]$ (4.13)
$={\bf u}^{n}+\Delta t\left\{\left(\frac{1}{6}+\frac{7}{2}C\right)\mathcal{L}({\bf u}^{n})+\left(\frac{2}{3}-10C\right)\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{6}+\frac{13}{2}C\right)\mathcal{L}\left[({\bf u}^{n+1})^{(k)}\right]\right\}$
$\quad\quad\;\;\;+C\Delta t^{2}\left\{\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{3}{2}\frac{\partial}{\partial t}\mathcal{L}\left[({\bf u}^{n+1})^{(k)}\right]\right\}-({\bf u}^{n+1})^{(k)}.$
5 Numerical tests
This section will validate the convergence order, stability, and competitiveness of the proposed two-stage fourth-order implicit (e.g., Implicit TSFO) scheme (3.35)-(3.36) by applying it to several classic stiff problems, with its fourth-order accuracy being verified through the extrapolation algorithm below,
$\text{Order}=\frac{\text{ln}[\text{Error}(\Delta t_{1})/\text{Error}(\Delta t_{2})]}{\text{ln}(\Delta t_{1}/\Delta t_{2})},$ (5.1)
especially, when $\Delta t_{1}=\Delta t$, $\Delta t_{2}=\Delta t/2$, we have
$\text{Order}_{L^{2}}(\Delta t)=\frac{\text{ln}[\text{Error}_{L^{2}}(\Delta t)/\text{Error}_{L^{2}}(\Delta t/2)]}{\text{ln}2},$ (5.2)
$\text{Order}_{L^{\infty}}(\Delta t)=\frac{\text{ln}[\text{Error}_{L^{\infty}}(\Delta t)/\text{Error}_{L^{\infty}}(\Delta t/2)]}{\text{ln}2},$ (5.3)
where
$\text{Error}_{L^{2}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{e}\lVert_{L^{2}},\quad\text{Error}_{L^{2}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{c}\lVert_{L^{2}},$ (5.4)
$\text{Error}_{L^{\infty}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{e}\lVert_{L^{\infty}},\quad\text{Error}_{L^{\infty}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{c}\lVert_{L^{\infty}},$ (5.5)
${\bf u}_{n}(\Delta t)$ is the numerical solution, ${\bf u}_{e}$ is the exact solution, and ${\bf u}_{c}$ is the reference solution generated by the classical fourth-order explicit Runge-Kutta method with a highly refined time step of $1.0\times 10^{-6}$. Additionally, a direct comparison is conducted between the proposed implicit TSFO scheme and the classical fourth-order implicit Runge-Kutta (e.g., Implicit RK4) method, focusing on stable time step size and convergence error, to verify the superiority of the former.
5.1 Linear System with Widely Separated Eigenvalues
This simple linear system,
$\begin{cases}\frac{du_{1}}{dt}=-1000u_{1}+1,\\ \frac{du_{2}}{dt}=-u_{2}+1,\end{cases}$ (5.6)
provides the most intuitive illustration of stiffness, with initial conditions: $u_{1}(0)=0,u_{2}(0)=0$ and time interval $t\in[0,10]$. The exact solution of this problem is
$\begin{cases}u_{1}(t)=0.001(1-e^{-1000t}),\\ u_{2}(t)=1-e^{-t}.\end{cases}$ (5.7)
The Jacobian matrix has eigenvalues at 1000-1000 and 1-1, representing two processes that decay at rates differing by a factor of one thousand. The component $u_{1}$ reaches its steady state almost instantaneously (by $t\approx 0.005$), while $u_{2}$ evolves slowly towards its steady state over a much longer period (until $t\approx 10$). For an explicit method, the step size is limited by the fastest process $(u_{1})$, requiring steps on the order of $1/1000$ even during the long interval where only the slow evolution of $u_{2}$ is of interest. Implicit methods excel here because their superior stability allows them to take large steps controlled by the accuracy requirement for $u_{2}$, efficiently skipping over the fast transient that has already decayed.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
      $\tau$=1.000000E+00 6.697969115862E-08 –– 8.869206919424E-04 ––
    $\tau/2$=5.000000E-01 4.145569039958E-09 4.0140816724 6.187833932367E-04 0.5193706265
    $\tau/4$=2.500000E-01 2.584340652531E-10 4.0037019911 1.466069642268E-04 2.0774808416
    $\tau/8$=1.250000E-01 1.614125216792E-11 4.0009718441 4.619753432439E-07 8.3099220681
  $\tau/16$=6.250000E-02 1.008647055151E-12 4.0002591596 9.623737490873E-12 ************
  $\tau/32$=3.125000E-02 6.303729887940E-14 4.0000719087 6.013786673943E-13 4.0002516928
  $\tau/64$=1.562500E-02 3.939772721240E-15 4.0000214067 3.758452794720E-14 4.0000629032
$\tau/128$=7.812500E-03 2.462345918528E-16 4.0000070497 2.349007393733E-15 4.0000157245
$\tau/256$=3.906250E-03 1.538963415636E-17 4.0000026093 1.468125620726E-16 4.0000039311
$\tau/512$=1.953125E-03 9.618514177224E-19 4.0000010755 9.175778879022E-18 4.0000009828
Table 1: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Linear Systems with Widely Separated Eigenvalues".
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
      $\tau$=1.000000E+00 6.697969115862E-08 –– 8.869204367202E-04 ––
    $\tau/2$=5.000000E-01 4.145569039958E-09 4.0140816724 6.187833919430E-04 0.5193702144
    $\tau/4$=2.500000E-01 2.584340652531E-10 4.0037019911 1.466069642060E-04 2.0774808388
    $\tau/8$=1.250000E-01 1.614125216792E-11 4.0009718441 4.619753175467E-07 8.3099221482
  $\tau/16$=6.250000E-02 1.008647055151E-12 4.0002591596 9.623737490765E-12 ************
  $\tau/32$=3.125000E-02 6.303729887940E-14 4.0000719087 6.013786673943E-13 4.0002516928
  $\tau/64$=1.562500E-02 3.939772721240E-15 4.0000214067 3.758452794720E-14 4.0000629032
$\tau/128$=7.812500E-03 2.462345918528E-16 4.0000070497 2.349007393733E-15 4.0000157245
$\tau/256$=3.906250E-03 1.538963415636E-17 4.0000026093 1.468125620726E-16 4.0000039311
$\tau/512$=1.953125E-03 9.618514177224E-19 4.0000010755 9.175778879022E-18 4.0000009828
Table 2: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Linear Systems with Widely Separated Eigenvalues".
5.2 Robertson’s Chemical Kinetics Problem
The Robertson problem is a classical stiff system that models a three-component chemical reaction describing hydrogen-oxygen-nitrogen kinetics. This benchmark problem is widely used to test numerical methods for stiff ordinary differential equations. The system represents the concentrations of different chemical species and exhibits characteristics of extreme stiffness due to the presence of both fast and slow time scales, with a stiffness ratio of approximately $10^{6}$. The system is governed by the following equations:
$\begin{cases}\frac{du_{1}}{dt}=-0.04u_{1}+10^{4}u_{2}u_{3},\\ \frac{du_{2}}{dt}=0.04u_{1}-10^{4}u_{2}u_{3}-3\times 10^{7}u_{2}^{2},\\ \frac{du_{3}}{dt}=3\times 10^{7}u_{2}^{2},\end{cases}$ (5.8)
with initial conditions $u_{1}(0)=1$, $u_{2}(0)=0$, $u_{3}(0)=0$. This problem is characterized by three distinct time scales: an ultrafast process ($\approx 10^{-8}$ s, governed by $k_{3}=3\times 10^{7}$), an intermediate process ($\approx 10^{-4}$ s, governed by $k_{2}=10^{4}$), and a slow process ($\approx 25$ s, governed by $k_{1}=0.04$). Therefore, to effectively validate the accuracy of numerical methods, a terminal time of $t_{\text{end}}=10$ s is chosen. This time point captures the complete transition from fast stiff dynamics to slow evolution, which tests the method’s capability to handle extreme stiffness transitions, while also avoiding the reduced sensitivity in error evaluation that typically occurs near equilibrium.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
        $\tau$=1.000000E-02 9.021072350964E-10 –– 3.854913505270E-09 ––
    $\tau/2$=5.000000E-03 2.580108817658E-10 1.8058670259 7.197422436148E-10 2.4211462484
    $\tau/4$=2.500000E-03 7.867372614174E-12 5.0354061892 7.128430647411E-11 3.3358239349
    $\tau/8$=1.250000E-03 2.330574682235E-12 1.7551961699 4.221448735445E-13 7.3997024919
  $\tau/16$=6.250000E-04 4.684507049390E-14 5.6366446928 1.549576641782E-13 1.4458640822
  $\tau/32$=3.125000E-04 1.740469668267E-15 4.7503486640 1.087571543380E-14 3.8326918986
  $\tau/64$=1.562500E-04 8.972457729393E-17 4.2778296404 6.977734145232E-16 3.9622078670
$\tau/128$=7.812500E-05 5.094883117507E-18 4.1383822665 4.389746093103E-17 3.9905492308
$\tau/256$=3.906250E-05 3.031651023500E-19 4.0708734516 2.748123534058E-18 3.9976187324
$\tau/512$=1.953125E-05 1.851325455787E-20 4.0334732502 1.718639521320E-19 3.9991079765
Table 3: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Robertson’s Chemical Kinetics Problem".
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
        $\tau$=1.000000E-02 6.379113507692E-10 –– 2.725942952427E-09 ––
    $\tau/2$=5.000000E-03 1.824484536839E-10 1.8058670259 5.089547175647E-10 2.4211461618
    $\tau/4$=2.500000E-03 5.563292362660E-12 5.0354061892 5.040760743533E-11 3.3358239189
    $\tau/8$=1.250000E-03 1.648030284841E-12 1.7551961699 2.985132390872E-13 7.3997027526
  $\tau/16$=6.250000E-04 3.312577599763E-14 5.6366446928 1.095759488245E-13 1.4458637453
  $\tau/32$=3.125000E-04 1.230746538617E-15 4.7503486640 7.690596265044E-15 3.8326919039
  $\tau/64$=1.562500E-04 6.344736420581E-17 4.2778296405 4.934198254849E-16 3.9622078681
$\tau/128$=7.812500E-05 3.602768766448E-18 4.1383822669 3.104141983081E-17 3.9905492311
$\tau/256$=3.906250E-05 2.143785699906E-19 4.0708734580 1.943293632771E-18 3.9976187332
$\tau/512$=1.953125E-05 1.309136414687E-20 4.0334733543 1.215309709126E-19 3.9991079877
Table 4: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Robertson’s Chemical Kinetics Problem".
5.3 Ozone Decomposition Reaction Problem
The ozone decomposition reaction system is a typical stiff system, characterized by the simultaneous presence of extremely short-lived, rapidly changing radical intermediates (such as oxygen atoms) and slowly varying major species (such as ozone and oxygen), resulting in significant disparities in the kinetic rates of different components. This problem is described by the following eight-dimensional system:
$\begin{cases}\frac{du_{1}}{dt}=-1.71u_{1}+0.43u_{2}+8.32u_{3}+0.0007,\\ \frac{du_{2}}{dt}=1.71u_{1}-8.75u_{2},\\ \frac{du_{3}}{dt}=-10.03u_{3}+0.43u_{4}+0.035u_{5},\\ \frac{du_{4}}{dt}=8.32u_{2}+1.71u_{3}-1.12u_{4},\\ \frac{du_{5}}{dt}=-1.745u_{5}+0.43u_{6}+0.43u_{7},\\ \frac{du_{6}}{dt}=-280u_{6}u_{8}+0.69u_{4}+1.71u_{5}-0.43u_{6}+0.69u_{7},\\ \frac{du_{7}}{dt}=280u_{6}u_{8}-1.81u_{7},\\ \frac{du_{8}}{dt}=-280u_{6}u_{8}+1.81u_{7},\end{cases}$ (5.9)
with initial conditions: $u_{1}(0)=1,u_{2}(0)=0,u_{3}(0)=0,u_{4}(0)=0,u_{5}(0)=0,u_{6}(0)=0,u_{7}(0),u_{8}(0)=0.0057$. In the numerical testing of the ozone decomposition reaction, three critical time points were selected to comprehensively evaluate the performance of the numerical method across different kinetic regimes. These points at 1.0 s, 10.0 s, and 321.8122 s respectively assess the algorithm’s capability in handling stiffness during rapid transients, convergence accuracy in the fully developed system phase, and steady-state precision after long-term integration.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
       $\tau$=5.0000000E-01 7.304648134165E-05 –– 2.103486775801E-03 ––
    $\tau/2$=2.5000000E-01 4.368777006510E-06 4.0635134059 5.402253376233E-05 5.2830776313
    $\tau/4$=1.2500000E-01 2.757893576099E-07 3.9855907777 2.740485668104E-06 4.3010578117
    $\tau/8$=6.2500000E-02 1.726194186584E-08 3.9979001126 1.695651923233E-07 4.0145196352
  $\tau/16$=3.1250000E-02 1.079170495833E-09 3.9996000520 1.059477506743E-08 4.0004151855
  $\tau/32$=1.5625000E-02 6.745110142817E-11 3.9999369994 6.621554154568E-10 4.0000392748
  $\tau/64$=7.8125000E-03 4.215681636995E-12 4.0000041759 4.138452076597E-11 4.0000067177
$\tau/128$=3.9062500E-03 2.634780405708E-13 4.0000112892 2.586529864712E-12 4.0000014966
$\tau/256$=1.9531250E-03 1.646728646603E-14 4.0000079786 1.616580758824E-13 4.0000003629
$\tau/512$=9.7656250E-04 1.029202139129E-15 4.0000045767 1.010362911214E-14 4.0000000900
Table 5: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=1.0$.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
       $\tau$=5.0000000E-01 5.257460149977E-05 –– 1.660009216431E-03 ––
    $\tau/2$=2.5000000E-01 3.259751779961E-06 4.0115319928 2.959645560584E-05 5.8096230285
    $\tau/4$=1.2500000E-01 2.062223567606E-07 3.9824894616 1.927179143601E-06 3.9408618221
    $\tau/8$=6.2500000E-02 1.293209504014E-08 3.9951728248 1.228186682972E-07 3.9718889151
  $\tau/16$=3.1250000E-02 8.089656191792E-10 3.9987338152 7.708323927517E-09 3.9939688558
  $\tau/32$=1.5625000E-02 5.057178443494E-11 3.9996737996 4.822552904827E-10 3.9985482291
  $\tau/64$=7.8125000E-03 3.160918871717E-12 3.9999167726 3.014846794572E-11 3.9996404695
$\tau/128$=3.9062500E-03 1.975603318786E-13 3.9999788049 1.884396367645E-12 3.9999103293
$\tau/256$=1.9531250E-03 1.234756723741E-14 3.9999945675 1.177766019856E-13 3.9999775955
$\tau/512$=9.7656250E-04 7.717237103705E-16 3.9999985829 7.361066198361E-15 3.9999943997
Table 6: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=1.0$.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
       $\tau$=5.0000000E-01 1.675891248573E-06 –– 8.943303121509E-06 ––
    $\tau/2$=2.5000000E-01 3.642515346050E-07 2.2019215776 1.390690585893E-06 2.6850063041
    $\tau/4$=1.2500000E-01 1.746476074086E-08 4.3824162667 8.930627055194E-08 3.9608961845
    $\tau/8$=6.2500000E-02 1.088293464093E-09 4.0043073341 5.748372386454E-09 3.9575360469
  $\tau/16$=3.1250000E-02 6.835681226013E-11 3.9928387090 3.619035911141E-10 3.9894761958
  $\tau/32$=1.5625000E-02 4.283237534322E-12 3.9963115285 2.265979384053E-11 3.9973987828
  $\tau/64$=7.8125000E-03 2.680473612197E-13 3.9981418489 1.416873790744E-12 3.9993515758
$\tau/128$=3.9062500E-03 1.676383727647E-14 3.9990636048 8.856455551941E-14 3.9998380121
$\tau/256$=1.9531250E-03 1.048081474569E-15 3.9995296457 5.535440071186E-15 3.9999595104
$\tau/512$=9.7656250E-04 6.551579640740E-17 3.9997642671 3.459674317438E-16 3.9999898781
Table 7: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=10.0$.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
       $\tau$=5.0000000E-01 1.627111303401E-06 –– 5.211344689802E-06 ––
    $\tau/2$=2.5000000E-01 3.536219744673E-07 2.2020331094 1.349971096619E-06 1.9487271620
    $\tau/4$=1.2500000E-01 1.695488884906E-08 4.3824346966 8.669408192396E-08 3.9608511959
    $\tau/8$=6.2500000E-02 1.056521292583E-09 4.0043075785 5.580256084178E-09 3.9575302772
  $\tau/16$=3.1250000E-02 6.636118877321E-11 3.9928383022 3.513197338642E-10 3.9894748094
  $\tau/32$=1.5625000E-02 4.158192376043E-12 3.9963113203 2.199711418803E-11 3.9973984426
  $\tau/64$=7.8125000E-03 2.602219852764E-13 3.9981417443 1.375437773761E-12 3.9993514911
$\tau/128$=3.9062500E-03 1.627443432415E-14 3.9990635523 8.597451492040E-14 3.9998379910
$\tau/256$=1.9531250E-03 1.017483834694E-15 3.9995296193 5.373558010232E-15 3.9999595051
$\tau/512$=9.7656250E-04 6.360313199880E-17 3.9997642539 3.358497322562E-16 3.9999898767
Table 8: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=10.0$.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
       $\tau$=5.0048554E-01 5.844310485027E-07 –– 3.774174108398E-08 ––
    $\tau/2$=2.5004833E-01 1.231981946167E-07 2.2435357849 4.679736139609E-07 ************
    $\tau/4$=1.2502416E-01 5.899164826275E-09 4.3843265848 3.004562111915E-08 3.9612005363
    $\tau/8$=6.2512082E-02 3.675895285249E-10 4.0043431451 1.934439284412E-09 3.9571673931
  $\tau/16$=3.1253006E-02 2.308023317497E-11 3.9928065125 1.217479412149E-10 3.9893871934
  $\tau/32$=1.5625744E-02 1.445943072441E-12 3.9962951902 7.621612240772E-12 3.9973773289
  $\tau/64$=7.8126824E-03 9.047964046463E-14 3.9981337125 4.765203882270E-13 3.9993462870
$\tau/128$=3.9062938E-03 5.658390449927E-15 3.9990595278 2.978444948245E-14 3.9998366978
$\tau/256$=1.9531350E-03 3.537566339472E-16 3.9995276037 1.861535568270E-15 3.9999591827
$\tau/512$=9.7656455E-04 2.211314986480E-17 3.9997632452 1.163453836558E-16 3.9999897963
Table 9: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=321.8122$.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
       $\tau$=5.0048554E-01 5.329967588582E-07 –– 3.531284647674E-08 ––
    $\tau/2$=2.5004833E-01 1.123553013200E-07 2.2435426293 4.267215439844E-07 ************
    $\tau/4$=1.2502416E-01 5.379973849055E-09 4.3843252272 2.739727677181E-08 3.9611905453
    $\tau/8$=6.2512082E-02 3.352376459368E-10 4.0043430876 1.763937593967E-09 3.9571610736
  $\tau/16$=3.1253006E-02 2.104892124693E-11 3.9928065125 1.110171783565E-10 3.9893857120
  $\tau/32$=1.5625744E-02 1.318684330103E-12 3.9962951920 6.949851378237E-12 3.9973769682
  $\tau/64$=7.8126824E-03 8.251644637160E-14 3.9981337135 4.345204108855E-13 3.9993461974
$\tau/128$=3.9062938E-03 5.160390443115E-15 3.9990595283 2.715928163278E-14 3.9998366754
$\tau/256$=1.9531350E-03 3.226221959847E-16 3.9995276039 1.697461925345E-15 3.9999591771
$\tau/512$=9.7656455E-04 2.016695175240E-17 3.9997632453 1.060908330213E-16 3.9999897949
Table 10: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=321.8122$.
5.4 Van der Pol Oscillator Problem
The van der Pol oscillator is a fundamental model in nonlinear dynamics, recognized for its characteristic stable limit cycle, an isolated periodic orbit that attracts neighboring trajectories. Governed by the nonlinear damping term $\nu(1-u^{2})\frac{du}{dt}$, the system exhibits self-excited oscillations whose behavior is critically influenced by the parameter $\nu$. At small values ($\nu<1$), the oscillator displays quasi-harmonic motion; as $\nu$ increases into the intermediate range ($1\leq\nu\leq 10$), distinct fast and slow time scales begin to emerge. For large $\nu$ values such as $\nu>10$, the system demonstrates pronounced stiffness, evolving into a relaxation oscillator characterized by slow energy accumulation followed by rapid discharge. In this study, we examine the strongly stiff regime with $\nu=100$, where the system exhibits extreme stiffness ($\nu\gg 100$). The model is formally expressed as the second-order equation
$\frac{d^{2}u}{dt^{2}}-\nu(1-u^{2})\frac{du}{dt}+u=0,$ (5.10)
which is commonly rewritten as a first-order system for numerical treatment by introducing $u_{1}=u$ and $u_{2}=\frac{du}{dt}$, yielding
$\begin{cases}\frac{du_{1}}{dt}=u_{2},\\ \frac{du_{2}}{dt}=\nu(1-u_{1}^{2})u_{2}-u_{1},\end{cases}$ (5.11)
with typical initial conditions $u_{1}(0)=2$, $u_{2}(0)=0$ over the time interval $t\in[0,100]$.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
        $\tau$=1.000000E-02 4.239641324385E-03 –– 3.224176703077E-03 ––
    $\tau/2$=5.000000E-03 1.333897566871E-05 8.3121505242 2.278940261915E-04 3.8224958012
    $\tau/4$=2.500000E-03 1.687551993730E-06 2.9826440252 1.574642176302E-05 3.8552671765
    $\tau/8$=1.250000E-03 1.061075442398E-07 3.9913328120 1.006550615658E-06 3.9675323991
  $\tau/16$=6.250000E-04 6.641293212617E-09 3.9979192304 6.326789505003E-08 3.9918023140
  $\tau/32$=3.125000E-04 4.152332733106E-10 3.9994702359 3.959878774008E-09 3.9979454275
  $\tau/64$=1.562500E-04 2.595462410976E-11 3.9998585548 2.475806070187E-10 3.9994860471
$\tau/128$=7.812500E-05 1.622207101622E-12 3.9999616736 1.547516403152E-11 3.9998717060
$\tau/256$=3.906250E-05 1.013864965219E-13 4.0000205949 9.672169552167E-13 3.9999713563
$\tau/512$=1.953125E-05 6.334338803941E-15 4.0005276704 6.044905950897E-14 4.0000477364
Table 11: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Van der Pol Oscillator Problem".
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
        $\tau$=1.000000E-02 4.239531568484E-03 –– 3.224093052535E-03 ––
    $\tau/2$=5.000000E-03 1.333862718107E-05 8.3121508669 2.278880751083E-04 3.8224960444
    $\tau/4$=2.500000E-03 1.687507904449E-06 2.9826440262 1.574601038175E-05 3.8552671938
    $\tau/8$=1.250000E-03 1.061047720459E-07 3.9913328121 1.006524318300E-06 3.9675324003
  $\tau/16$=6.250000E-04 6.641119700385E-09 3.9979192304 6.326624209600E-08 3.9918023141
  $\tau/32$=3.125000E-04 4.152224248130E-10 3.9994702359 3.959775317138E-09 3.9979454275
  $\tau/64$=1.562500E-04 2.595394601219E-11 3.9998585548 2.475741386603E-10 3.9994860471
$\tau/128$=7.812500E-05 1.622164719398E-12 3.9999616736 1.547475972317E-11 3.9998717060
$\tau/256$=3.906250E-05 1.013838476707E-13 4.0000205949 9.671916854432E-13 3.9999713563
$\tau/512$=1.953125E-05 6.334173311281E-15 4.0005276704 6.044748020038E-14 4.0000477364
Table 12: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Van der Pol Oscillator Problem".
Numerical results confirm that the proposed two-step fourth-order implicit (Implicit TSFO) scheme effectively solves stiff problems using large time steps without compromising fourth-order accuracy. When compared to the classical implicit Runge-Kutta (Implicit RK4) method, our scheme reduces the numerical error by about an order of magnitude under identical time steps and offers superior stability that enables even larger step sizes.
6 Conclusions and future work
This paper presents a novel two-stage fourth-order implicit scheme. We rigorously establish its fourth-order accuracy through mathematical analysis and systematically examine the scheme’s stability, deriving the conditions under which A-stability can be achieved. A detailed implementation procedure based on Newton iteration is provided. The effectiveness of the proposed method is validated through numerical experiments on several classical strongly stiff benchmark cases. Comparisons with classical explicit and implicit fourth-order Runge-Kutta methods demonstrate that the proposed scheme possesses superior stability regions and smaller numerical errors.
In future work, we will combine this implicit scheme with its explicit counterpart—a spatiotemporally coupled two-stage fourth-order method—to solve fluid dynamics problems characterized by strongly stiff source terms, such as those encountered in chemical reactions. Compared to other types of implicit schemes, the proposed method incorporates temporal derivatives of physical quantities, which is expected to enhance both computational accuracy and compactness of the scheme.
7 Acknowledgments
Zhixin Huo’s research work has been supported by the Key Program of Henan Higher Education Institutions (Grant No. 26A110007), the Young Talents Fund of Henan Province (Grant No. 252300423500), the "Double First-Class" Project of the School of Geomatics, Henan Polytechnic University (Grant No. BSJJ202306), and the Doctoral Startup Foundation of Henan Polytechnic University (Grant No. B2024-60).

### v2

@@
-A Two-Stage Fourth-Order Implicit Scheme for Stiff problems
+An Implicit Two-Stage Fourth-Order Temporal Discretization Scheme of Lax-Wendroff Type for Hydrodynamic Problems with Stiff Source Terms Part I: Formulation, Stability Analysis, and Newton’s Iteration
 Abstract
-This paper presents a novel two-stage fourth-order implicit scheme designed to overcome the limitations of existing spatiotemporal coupling two-stage fourth-order methods, which have thus far been confined to explicit frameworks.
-In such frameworks, computational efficiency is severely constrained by the CFL condition when addressing stiff problems, and they fundamentally conflicts with spatiotemporal coupled methods whose core advantage lies in embedding stiff source terms.
-By fully leveraging temporal derivatives of physical variables within a rigorously derived compact mathematical framework, the scheme achieves fourth-order temporal accuracy in only two stages.
-Furthermore, a sufficient condition for A-stability is established through systematic theoretical and numerical investigation, and a Newton iterative procedure is provided to accelerate convergence.
-Extensive numerical experiments on classical stiff problems confirm the method’s effectiveness and competitiveness.
-Key words: two-stage fourth-order; temporal derivatives; implicit scheme; A-stability; stiff problems; Newton iteration
+Existing two-stage fourth-order (TSFO) methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] are confined to explicit frameworks and can only be applied to homogeneous hydrodynamic problems describing idealized states.
+More importantly, the ability of these methods to achieve fourth-order time accuracy in merely two stages relies on utilizing temporal derivatives of physical quantities, which are obtained through generalized Riemann problem (GRP) solvers.
+While the advantage of the GRP solvers [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] over traditional Riemann solvers [28, 29] lies precisely in their ability to naturally incorporate the effects of source terms.
+Specifically, by employing the Lax-Wendroff methodology, the temporal derivatives of physical quantities are converted into spatial derivatives of fluxes and source terms through the governing equations, thereby avoiding errors arising from operator splitting.
+Therefore, to extend the applicability of TSFO methods to hydrodynamic problems with stiff source terms that describe real physical phenomena, and to fully leverage the advantages of these GRP solvers [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], there is an urgent need to extending TSFO methods from explicit frameworks to implicit ones.
+The present work lays a solid theoretical foundation and provides a concrete implementation pathway for this research direction.
+Firstly, this paper derives, through rigorous mathematical analysis, an implicit temporal discretization scheme that achieves fourth-order accuracy in just two stages.
+Subsequently, sufficient conditions for A‑stability of the proposed implicit scheme are established via systematic theoretical analysis, and a corresponding Newton iteration procedure is formulated to accelerate convergence.
+Finally, the proposed implicit scheme is extensively validated using classical stiff benchmark problems, confirming its effectiveness and competitiveness.
+Numerical results demonstrate that the proposed scheme indeed achieves fourth-order time accuracy in two stages and, compared to the classical fourth‑order implicit Runge‑Kutta method, offers a larger stable time step and smaller convergence error.
+Key words: two-stage fourth-order (TSFO); generalized Riemann problem (GRP); Lax-Wendroff methodology; implicit scheme; stiff source term; A-stability
 1 Introduction
-The generalized Riemann problem (GRP) fundamentally differs from the classical Riemann problem through its use of spatially varying initial data, which generates non-self-similar solutions with non-zero time derivatives.
-This critical advancement enables the natural incorporation of physical source terms and other essential effects (e.g., continuous entropy variation) into the temporal evolution, establishing GRP as an indispensable framework for capturing complex physical mechanisms.
-The simplest GRP form replaces piecewise constant with piecewise linear initial data in homogeneous systems, first explored by Kolgan [1] and further developed by van Leer [2, 3].
-Toro and Titarev [4] later extended GRP to include source terms using the Lax-Wendroff procedure, forming the basis of arbitrarily high-order (ADER) schemes [5, 6] applicable to hydrodynamics, magnetohydrodynamics [7], reaction-diffusion [8], and stiff systems [9], though these linearized solvers struggle under extreme thermodynamic conditions.
-Ben-Artzi et al.
-[10, 11, 12] were the first to develop the GRP solver for compressible fluids, which can be considered as the nonlinear version of the ADER solver.
-By incorporating thermodynamic effects, the GRP solver can accurately resolve strong waves, making it valuable under extreme conditions [13].
-The GRP solver has since been successfully applied to duct flow [14], reactive flow [15], shallow water equations [16, 17], multi-material and multiphase flow [18, 19, 20, 21], magnetohydrodynamics [22], relativistic hydrodynamics [23, 24, 25], blood flow modeling [26], and radiation hydrodynamics [27].
-The inherent spatiotemporal coupling in fluid dynamics necessitates numerical algorithms that preserve such correlations, which conventional Riemann solvers fail to achieve as they only provide piecewise constant states for constructing instantaneous Cauchy fluxes.
-This limitation leads to error accumulation over time in flows with strong discontinuities.
-In contrast, generalized Riemann solvers (e.g., ADER or GRP) provide temporal derivatives of physical quantities, enabling construction of time-interval fluxes with inherent spatiotemporal coupling.
-These fluxes exhibit Lipschitz continuity with respect to boundary perturbations, allowing temporal effects to enhance spatial regularity.
-Consequently, errors from strong discontinuities can be directly controlled through the time step, enabling accurate long-term resolution while establishing consistency between mathematical weak solutions and physical integral balance laws [28, 29].
-Runge-Kutta methods, originally developed by Carl Runge [30] in 1895 and Martin Wilhelm Kutta [31] in 1901, constitute a family of classical numerical techniques for solving initial-value problems of ordinary differential equations that comprise both explicit and implicit forms.
-The explicit variants, exemplified by the classical fourth-order RK4 scheme, achieve high-order accuracy through weighted averages of function values at multiple stages, offering implementation simplicity without requiring calculation of higher-order derivatives.
-However, their stability remains constrained by the CFL condition, rendering them inefficient for stiff problems.
-In contrast, Implicit Runge-Kutta methods proposed by Kuntzmann [32] and Butcher [33], such as Radau and Gauss-type schemes, compute slopes at future time steps by solving systems of nonlinear equations.
-Although computationally more expensive, they exhibit superior numerical stability, making them particularly suitable for stiff differential equations.
-Recently, Li et al.
-developed a two-stage fourth-order time advancement method [34], which makes full use of the temporal derivatives of physical quantities, enabling the scheme to achieve fourth-order time accuracy in only two stages.
-This method successfully combines the simplicity advantages of Runge-Kutta-type methods with the spatiotemporally coupling properties of the GRP solver.
-Then they details the characteristics of this class of methods [35]: (i) Computational efficiency: Under the same conditions, its computational efficiency is twice that of the Runge-Kutta method, which requires four stages to achieve fourth-order time accuracy [36]; (ii) Compactness: Since the number of time advancement steps is reduced by half, the stencil size is also halved; the spatial-temporal coupled HWENO-type reconstruction [37, 38] can make the computational scheme more compact; (iii) Stability: It has been proven that its stability region is larger than that of Runge-Kutta methods [39]; (iv) Compatibility: The method can be easily combined with other methods, such as the GKS solver [40], the multi-moment method [41], and the CRP reconstruction technique [42].
-However, despite the aforementioned advantages, current spatiotemporally coupled two-step fourth-order methods still suffer from significant limitations.
-On the one hand, such methods are currently confined to explicit frameworks.
-As a result, their time step is severely constrained by the CFL stability condition, and this restriction in turn drastically reduces their computational efficiency when dealing with stiff problems.
-On the another hand, the stiffness of a problem is often predominantly manifested in its source terms.
-For instance, in problems involving multi-scale physics like combustion or chemically reacting flows, the source terms can introduce significant stiffness.
-The key advantage of the GRP solver over the conventional Riemann solver lies in its ability to provide the temporal derivatives of physical quantities, and this derivative information inherently encodes the effects of the source terms from the governing equations.
-However, a critical paradox emerges: this same temporal derivative data is utilized in two distinct ways: it is essential for constructing the time-interval fluxes and serves as the cornerstone for building the two-stage fourth-order explicit scheme.
-The former (time-interval fluxes) can successfully incorporate information from stiff source terms, whereas the latter (the explicit scheme itself) is fundamentally unsuitable for solving stiff problems due to its explicit nature.
-This inherent incompatibility points to a significant limitation in the current methodological framework.
-Inaddition, the incorporation of temporal derivatives may allows the proposed method to achieve not only higher computational accuracy but also a more compact formulation compared to existing implicit schemes.
-Based on the above considerations, we conclude that there is an urgent need to develop two-stage fourth-order implicit methods to address multi-scale problems such as those involving strong stiffness.
-As the foundational component of this research on spatiotemporally coupled two-stage fourth-order implicit schemes, hereafter referred to as Implicit TSFO, we develop an implicit framework that attains fourth-order accuracy in two stages through the incorporation of temporal derivatives of physical quantities.
-Section 2 presents the detailed formulation of the scheme via rigorous mathematical derivation.
-This is followed in Section 3 by a thorough analysis of the stability conditions of the implicit method.
-Section 4 outlines the implementation procedure based on Newton iteration.
-In Section 5, the efficacy of the proposed scheme is demonstrated through several canonical numerical examples, supplemented by comparative studies with classical fourth-order implicit and explicit Runge-Kutta methods.
-Finally, Section 6 concludes the paper and suggests potential directions for future research.
-2 The Construction of the Implicit TSFO Scheme
+Hydrodynamic problems with stiff source terms establish a fundamental mathematical framework for describing complex physical phenomena in the real world.
+This framework distinguishes physical processes according to their characteristic time scales: the hydrodynamic equations govern slow, wave-dominated macroscopic dynamics, while stiff source terms capture fast, localized microscopic interactions.
+Classical examples include shock-reaction coupling in detonation waves, pollutant-photochemical reactions in atmospheric transport, shock-radiation coupling in astrophysical flows, and the co-evolution of magnetohydrodynamic waves with atomic-scale processes in fusion plasmas.
+By contrast, homogeneous hydrodynamic equations reduce such complex flow systems in real physical environments to purely convective systems characterized by a single time scale.
+In numerical implementation, the latter, which involves only one time scale, can typically be solved efficiently using explicit schemes.
+The former, however, contains multiple distinct fast and slow time scales within the same system.
+If a purely explicit scheme is applied, the time step constrained by the CFL stability condition will be limited by the smallest time scale, leading to a sharp decline in computational efficiency.
+Therefore, such systems often require implicit or semi-implicit schemes, in which slowly varying fluid convection is treated explicitly while rapidly varying stiff source terms are handled implicitly, enabling efficient and stable numerical simulation across multiple scales.
+The Lax-Wendroff method, serving as a discrete implementation of the Cauchy-Kovalevskaya theorem, establishes and popularizes the fundamental paradigm of directly coupling temporal and spatial derivatives through the governing equations to achieve high-order discretization.
+It inherently supports the embedding of stiff source terms.
+Its core concept can be metaphorically described as predicting temporal evolution from spatial structures.
+Specifically, the method employs a Taylor expansion in time and, by means of the governing equations, transforms higher-order temporal derivatives into combinations of spatial derivatives that incorporate both flux terms and source terms.
+As a result, temporal evolution information from source terms is automatically integrated into the construction of local time-advancement operators, thereby enabling a unified description of multiscale physical processes and avoiding numerical errors introduced by operator-splitting methods that separately treat fluxes and sources.
+This characteristic makes the Lax-Wendroff framework particularly suitable for developing implicit high-order schemes, offering an effective mathematical framework for solving complex multiscale flow problems in real-world physics.
+The philosophical principle of Lax-Wendroff, which couples space and time through the governing equations themselves, has become a cornerstone for developing advanced numerical methods.
+One of its key applications is the construction of high‑accuracy generalized Riemann problem (GRP) solvers.
+For instance, Men’shov’s modified GRP solver [14] enhances the resolution of complex wave interactions; ADER solver [15, 16, 17, 18, 19, 20, 21] enables arbitrary‑order accuracy in both space and time; and Ben-Artiz’s nonlinear GRP solver [22, 23, 24, 25, 26, 27] provides precise treatment of strong discontinuities in compressible flows.
+These modern descendants are capable of handling hydrodynamic systems with stiff source terms within a unified, unsplit framework, thereby avoiding operator‑splitting errors and preserving the asymptotic properties of complex multiscale physical systems.
+Furthermore, such GRP solvers are essential for simulating real‑world flow fields containing shocks, turbulence, material interfaces, and intricate physical phenomena, where the regularity of the flow is typically very weak.
+This stems from the following reasons.
+Classical Riemann solvers like HLL solver [28] and Roe solver [29] are limited to resolving instantaneous states to construct Cauchy-type instantaneous fluxes, often leading to accumulated numerical errors near strong discontinuities.
+In contrast, GRP solvers incorporate temporal derivative information to capture local evolutionary trends, enabling construction of temporal-interval fluxes.
+According to finite-volume principles, these temporal-interval fluxes compensate for spatial regularity deficiencies through their built-in temporal evolution, allowing accurate resolution of strong discontinuities [30, 31].
+The philosophy of the Lax-Wendroff method is applicable not only to the development of high‑accuracy GRP solvers, but also to the construction of high‑order numerical schemes.
+The classical Lax‑Wendroff procedure can achieve arbitrarily high‑order temporal accuracy in a single step, but it requires repeated differentiation of the governing equations.
+In contrast, Runge‑Kutta methods [32, 33, 34, 35] achieve high‑order temporal accuracy without differentiating the governing equations, yet they involve multi‑stage time integration, which inevitably enlarges the computational stencil.
+To effectively balance these two strategies, Li et al.
+[1] proposed a two‑stage fourth‑order (TSFO) temporal discretization scheme that makes full use of the temporal derivative information of physical quantities and attains fourth‑order accuracy in only two stages.
+The research team later systematically elaborated the advantages of such kind of methods [3]: (1) Computational efficiency: Under comparable conditions, its computational efficiency is twice that of a four‑stage, fourth-order Runge‑Kutta method [4]; (2) Compactness: Because the number of time‑advancement stages is halved, the size of the computational stencil is correspondingly reduced; GRP-solver-based spatiotemporally coupled HWENO‑type reconstruction techniques [5, 6] further enhance the compactness of the scheme; (3) Stability: Its stability region has been proven to be larger than that of Runge‑Kutta methods [7]; (4) Compatibility: The method can be flexibly combined with various other numerical techniques, such as the GKS solver [8], the multi‑moment method [10], and the CRP reconstruction technique [11].
+Nevertheless, despite these advantages, existing TSFO methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] remain confined to explicit frameworks.
+As a result, they are only applicable to homogeneous hydrodynamic systems and cannot effectively address problems involving stiff source terms.
+What is more important, the capability of such schemes to achieve fourth-order temporal accuracy in merely two stages originates from their design, which fully exploits temporal derivative information of the physical quantities, which are derived from Lax-Wendroff-based GRP solvers.
+And the key advantage of these solvers lies precisely in their ability to integrate spatial derivatives of fluxes and source terms simultaneously into the time derivatives of physical quantities through the governing equations, all within a unified framework, thereby eliminating numerical errors that typically arise from operator splitting.
+Consequently, there is a clear and pressing need to develop an implicit TSFO temporal discretization scheme.
+The development of such a scheme would not only extend the applicability of TSFO methods from homogeneous hydrodynamic problems to systems with stiff source terms, but would also fully harness the Lax-Wendroff methodology’s inherent capacity for seamless stiff-source integration.
+As the theoretical foundation and innovative core of this study, the paper is structured as follows.
+Chapter 2 introduces a rigorously derived implicit temporal discretization scheme that achieves fourth-order accuracy in only two stages by fully utilizing the temporal derivative information of the physical quantities.
+Chapter 3 establishes sufficient conditions for the A-stability of the proposed scheme through systematic theoretical analysis.
+Chapter 4 further develops a Newton‑iteration formulation of the scheme to accelerate convergence.
+Chapter 5 validates the accuracy and stability of the method using a series of classical stiff benchmark problems, and confirms its effectiveness and competitiveness by comparing its performance with classical implicit and explicit Runge‑Kutta schemes.
+Finally, Chapter 6 summarizes the work and outlines potential directions for future research.
+2 Formulation for the Implicit TSFO Scheme
 Consider the time-dependent equations
 $\frac{\partial\bf u}{\partial t}=\mathcal{L}(\bf u),$ (2.1)
-where we assume that the operator $\mathcal{L}$ is sufficiently smooth, meaning all required derivatives exist. We write (2.1) as
+where we assume that the operator $\mathcal{L}$ is sufficiently smooth, meaning all required derivatives exist.
+In the follows, a two-stage fourth-order implicit scheme for system (2.1) is derived through rigorous mathematical analysis. (2.1) is rewritten as
 ${\bf u}^{n+\theta}={\bf u}^{n}+\int_{t_{n}}^{t_{n}+\theta\Delta t}\mathcal{L}[{\bf u}(t)]dt,$ (2.2)
-where ${\bf u}^{n+\theta}={\bf u}(t_{n}+\theta\Delta t)$, $\theta\in[0,1]$, denote $\mathcal{Q}(t)=\mathcal{L}[{\bf u}(t)]$, then
+where ${\bf u}^{n+\theta}={\bf u}(t_{n}+\theta\Delta t)$, $\theta\in[0,1]$. Denote $\mathcal{Q}(t)=\mathcal{L}[{\bf u}(t)]$, then
 $\mathcal{Q}^{\prime}(t)=\mathcal{L}_{\bf u}\mathcal{L},\quad\mathcal{Q}^{\prime\prime}(t)=\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L},\quad\mathcal{Q}^{\prime\prime\prime}(t)=\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L},$ (2.3)
@@
 $\quad\quad\quad\quad\;\;\;+\frac{(\Delta t)^{4}}{24}[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})+\mathcal{O}(\Delta t^{5}).$
-We construct a two-stage fourth-order implicit scheme by utilizing the values of physical quantities along with their temporal derivatives. The proposed scheme is formulated as follows:
+The proposed two-stage fourth-order implicit scheme is given as follows:
 Stage 1.
-We construct the following implicit scheme by utilizing the physical quantities at the time level $t_{n}$ and $t_{n+\frac{1}{2}}$ to obtain the solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$,
+The solution at the intermediate time level, $\mathbf{u}^{n+\frac{1}{2}}$, is obtained by utilizing the physical quantities and their temporal derivatives at the time levels $t_{n}$ and $t_{n+\frac{1}{2}}$,
 ${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\Delta t\left[A_{1}\mathcal{L}({\bf u}^{n})+A_{2}\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]+\Delta t^{2}\left[B_{1}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+B_{2}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right],$ (2.7)
-where the coefficients $A_{1}$, $A_{2}$, $B_{1}$, $B_{2}$ are to be determined to achieve fourth-order accuracy in Stage 1.
+where the coefficients $A_{1}$, $A_{2}$, $B_{1}$, $B_{2}$ are to be determined to achieve fourth-order accuracy.
 Stage 2.
-We construct the following implicit scheme by utilizing the physical quantities at the time level $t_{n}$, $t_{n+\frac{1}{2}}$ and $t_{n+1}$ to obtain the solution at the next time level ${\bf u}^{n+1}$,
+The solution at the next time level, $\mathbf{u}^{n+1}$, is obtained by utilizing the physical quantities and their temporal derivatives at the time levels $t_{n}$, $t_{n+\frac{1}{2}}$ and $t_{n+1}$,
 ${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left[A_{3}\mathcal{L}({\bf u}^{n})+A_{4}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+A_{5}\mathcal{L}({\bf u}^{n+1})\right]$ (2.8)
 $+\Delta t^{2}\left[B_{3}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+B_{4}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+B_{5}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right],$
-where the coefficients $A_{3}$, $A_{4}$, $A_{5}$, $B_{3}$, $B_{4}$, $B_{5}$ are to be determined to achieve fourth-order accuracy in $\bf Stage2$.
+where the coefficients $A_{3}$, $A_{4}$, $A_{5}$, $B_{3}$, $B_{4}$, $B_{5}$ are to be determined to achieve fourth-order accuracy.
 By treating ${\bf u}^{n}$, ${\bf u}^{n+\frac{1}{2}}$, and ${\bf u}^{n+1}$ as the exact solution and using Taylor series expansions, we obtain
@@
 $=[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})+({\bf u}^{n+\theta}-{\bf u}^{n})[\mathcal{L}_{\bf uu}\mathcal{L}+\mathcal{L}_{\bf u}^{2}]({\bf u}^{n})$
-$\quad\quad\quad\quad\quad\quad+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2}[\mathcal{L}_{\bf uuu}\mathcal{L}+3\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}]+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{3}].$
+$\quad\quad\quad\quad\quad\;\;\;+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2}[\mathcal{L}_{\bf uuu}\mathcal{L}+3\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}]+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{3}].$
 Substituting (2.11), (2.13), (2.15) into (2.7), we obtain
 ${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\Delta t\left(A_{1}+A_{2}\right)\mathcal{L}({\bf u}^{n})$ (2.16)
-$\quad\quad\quad\quad\quad\;\;+(\Delta t)^{2}\left(\frac{1}{2}A_{2}+B_{1}+B_{2}\right)[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})$
-$\quad\quad\quad\quad\quad\;\;+(\Delta t)^{3}\left(\frac{1}{8}A_{2}+\frac{1}{2}B_{2}\right)[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$
-$\quad\quad\quad\quad\quad\;\;+(\Delta t)^{4}\left(\frac{1}{48}A_{2}+\frac{1}{8}B_{2}\right)[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})$
-$\quad\quad\quad\quad\quad\;\;+\mathcal{O}[(\Delta t)^{5}].$
-Compare (2.5) with (2.16), we obtain that the coefficients must satisfy the following equation relationship,
-$\begin{cases}A_{1}+A_{2}=\frac{1}{2},\\ \frac{1}{2}A_{2}+B_{1}+B_{2}=\frac{1}{8},\\ \frac{1}{8}A_{2}+\frac{1}{2}B_{2}=\frac{1}{48},\\ \frac{1}{48}A_{2}+\frac{1}{8}B_{2}=\frac{1}{384},\end{cases}$ (2.17)
-Solve the above system of equations, we obtain the unique solution
+$\quad\quad\quad\quad\;\;\;+(\Delta t)^{2}\left(\frac{1}{2}A_{2}+B_{1}+B_{2}\right)[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})$
+$\quad\quad\quad\quad\;\;\;+(\Delta t)^{3}\left(\frac{1}{8}A_{2}+\frac{1}{2}B_{2}\right)[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$
+$\quad\quad\quad\quad\;\;\;+(\Delta t)^{4}\left(\frac{1}{48}A_{2}+\frac{1}{8}B_{2}\right)[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})$
+$\quad\quad\quad\quad\;\;\;+\mathcal{O}[(\Delta t)^{5}].$
+Compare (2.5) with (2.16), we obtain that the coefficients must satisfy the following system of equations,
+$\begin{cases}A_{1}+A_{2}=\frac{1}{2},\\ \frac{1}{2}A_{2}+B_{1}+B_{2}=\frac{1}{8},\\ \frac{1}{8}A_{2}+\frac{1}{2}B_{2}=\frac{1}{48},\\ \frac{1}{48}A_{2}+\frac{1}{8}B_{2}=\frac{1}{384}.\end{cases}$ (2.17)
+Solve the above system of equations, we obtain the unique solution,
 $A_{1}=\frac{1}{4},\quad A_{2}=\frac{1}{4},\quad B_{1}=\frac{1}{48},\quad B_{2}=-\frac{1}{48}.$ (2.18)
@@
 $+\Delta t^{4}\left(\frac{1}{48}A_{4}+\frac{1}{6}A_{5}+\frac{1}{8}B_{4}+\frac{1}{2}B_{5}\right)[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})$
-$+\mathcal{O}(\Delta t^{5})$
+$+\mathcal{O}(\Delta t^{5}).$
 The comparison of (2.6) and (2.19), gives
-$\begin{cases}A_{3}+A_{4}+A_{5}=1,\\ \frac{1}{2}A_{4}+A_{5}+B_{3}+B_{4}+B_{5}=\frac{1}{2},\\ \frac{1}{8}A_{4}+\frac{1}{2}A_{5}+\frac{1}{2}B_{4}+B_{5}=\frac{1}{6},\\ \frac{1}{48}A_{4}+\frac{1}{6}A_{5}+\frac{1}{8}B_{4}+\frac{1}{2}B_{5}=\frac{1}{24},\end{cases}$ (2.20)
+$\begin{cases}A_{3}+A_{4}+A_{5}=1,\\ \frac{1}{2}A_{4}+A_{5}+B_{3}+B_{4}+B_{5}=\frac{1}{2},\\ \frac{1}{8}A_{4}+\frac{1}{2}A_{5}+\frac{1}{2}B_{4}+B_{5}=\frac{1}{6},\\ \frac{1}{48}A_{4}+\frac{1}{6}A_{5}+\frac{1}{8}B_{4}+\frac{1}{2}B_{5}=\frac{1}{24}.\end{cases}$ (2.20)
 Solve the above system of equations, we obtain
-$\begin{cases}A_{3}=\frac{1}{6}+4C+\frac{1}{2}D,\\ A_{4}=\frac{2}{3}-8C+2D,\\ A_{5}=\frac{1}{6}+4C-\frac{5}{2}D,\\ B_{3}=C,\\ B_{4}=D,\\ B_{5}=\frac{1}{2}D-C.\end{cases}$ (2.21)
+$\begin{cases}A_{3}=\frac{1}{6}+4C+\frac{1}{2}D,\\ A_{4}=\frac{2}{3}-8C+2D,\\ A_{5}=\frac{1}{6}+4C-\frac{5}{2}D,\\ B_{3}=C,\\ B_{4}=D,\\ B_{5}=\frac{1}{2}D-C,\end{cases}$ (2.21)
 where $C$, $D$ are real parameters to be determined.
-Substituting (2.18) and (2.21) into (2.7) and (2.8) respectively, we obtain the specific form of the desired two-stage fourth-order implicit (e.g., Implicit TSFO) scheme as follows:
+Substituting (2.18) and (2.21) into (2.7) and (2.8) respectively, we obtain the specific form of the desired implicit two-stage fourth-order (Implicit TSFO) temporal discretization scheme as follows:
 Stage 1. Iteratively solve the following implicit scheme to obtain the solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$,
@@
 ${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left[\left(\frac{1}{6}+4C+\frac{1}{2}D\right)\mathcal{L}({\bf u}^{n})+\left(\frac{2}{3}-8C+2D\right)\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{6}+4C-\frac{5}{2}D\right)\mathcal{L}({\bf u}^{n+1})\right]$ (2.23)
-$+\Delta t^{2}\left[C\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+D\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{2}D-C\right)\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right].$
-3 The Stability Analysis of the Implicit TSFO Scheme
+$+\Delta t^{2}\left[C\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+D\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{2}D-C\right)\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right],$
+where $C$, $D$ are real parameters to be determined.
+3 Stability Analysis for the Implicit TSFO Scheme
 To examine the stability of the implicit scheme (2.22)-(2.23), we consider the following model equation
@@
 ${\bf u}^{n+\frac{1}{2}}=R(z){\bf u}^{n},$ (3.5)
-where
+where the amplification factor
 $R(z)=\frac{1+\frac{z}{4}+\frac{z^{2}}{48}}{1-\frac{z}{4}+\frac{z^{2}}{48}}=1+\frac{24}{z}+\mathcal{O}\left(\frac{1}{z^{2}}\right).$ (3.6)
@@
 ${\bf u}^{n+1}=G(z){\bf u}^{n},$ (3.7)
-where
+where the amplification factor
 $G(z)$ $=\frac{1+z\left[\left(\frac{1}{6}+4C+\frac{1}{2}D\right)+\left(\frac{2}{3}-8C+2D\right)R(z)\right]+z^{2}\left[C+DR(z)\right]}{1-\left(\frac{1}{6}+4C-\frac{5}{2}D\right)z-\left(\frac{1}{2}D-C\right)z^{2}}$ (3.8)
@@
 Our objective in the following is to determine the range of $C$ for which the Implicit TSFO scheme (2.22)-(2.23) is A-stable.
-Achieving this requires not only need condition (3.9) above but also need the following condition:
+Achieving this requires not only need condition (3.9) but also need the following condition:
 $\sup_{\text{Re}(z)\leq 0}|G(z)|\leq 1.$ (3.12)
-The theoretical foundation for verifying A-stability is based on the maximum modulus principle: If $G(z)$ is analytic in the closed left half-plane $\text{Re}(z)\leq 0$, then the maximum of $\lvert G(z)\lvert$ occurs on the boundary, specifically on: the imaginary axis $z=iy,\;y\in\mathbb{R}$ and at infinity $z\to\infty$.
-Therefore, to achieve the A-stability of the Implicit TSFO scheme (2.22)-(2.23), we need to examine the following three conditions:
+The theoretical foundation for verifying the condition (3.12) is based on the Maximum Modulus Principle: If $G(z)$ is analytic in the closed left half-plane $\text{Re}(z)\leq 0$, then the maximum of $\lvert G(z)\lvert$ occurs on the boundary, specifically on: the imaginary axis $z=iy,\;y\in\mathbb{R}$ and at infinity $z\to\infty$.
+Therefore, to achieve the condition (3.12) for the Implicit TSFO scheme (2.22)-(2.23), we need to examine the following three conditions:
 Condition 1: $G(z)$ is analytic in the closed left half-plane, i.e.,
-$G\in\mathcal{H}(\mathbb{C}_{-}),$ (3.13)
+$G(z)\in\mathcal{H}(\mathbb{C}_{-}),$ (3.13)
 where $\mathbb{C}_{-}=\{z\in\mathbb{C}:\mathrm{Re}(z)\leq 0\}$.
-Condition 2: The $L_{\infty}$-norm of $G$ on the imaginary axis is bounded by 1, i.e.,
+Condition 2: The $L_{\infty}$-norm of $G(z)$ on the imaginary axis is bounded by 1, i.e.,
 $\|G(i\cdot)\|_{L_{\infty}(\mathbb{R})}\leq 1,$ (3.14)
@@
 $\sup_{y\in\mathbb{R}}|G(iy)|\leq 1.$ (3.15)
-Condition 3: $G$ is bounded at infinity with magnitude not exceeding 1, i.e.,
+Condition 3: $G(z)$ is bounded at infinity with magnitude not exceeding 1, i.e.,
 $|G(\infty)|\leq 1.$ (3.16)
@@
 Examination of Condition 1, i.e., (3.13).
-The first requirement for A-stability is that the amplification factor $G(z)$ must be analytic in the closed left half-plane $\mathbb{C}_{-}$.
-This is equivalent to requiring that all poles of $G(z)$ lie strictly in the right half-plane.
+Condition 1 requires the amplification factor $G(z)$ to be analytic in the closed left half-plane $\mathbb{C}_{-}$, which is equivalent to requiring that all poles of $G(z)$ lie strictly in the right half-plane.
 The denominator of $G(z)$ is given by:
@@
 ${\textnormal{Re}}(z)>0,$ (3.25)
-i.e., all roots of the equation (3.21) lie in the right half-plane, in another words, the amplification factor $G(z)$ is analytic in $\textnormal{Re}(z)\leq 0$.
+i.e., all roots of the equation (3.21) lie in the right half-plane, in another words, the amplification factor $G(z)$ is analytic in the closed left half-plane.
 Examination of Condition 2, i.e., (3.14)-(3.15).
@@
 $\psi(y)=2\text{arg}\left(1+i\frac{y}{4}-\frac{y^{2}}{48}\right)=2\text{arctan}\left(\frac{y/4}{1-y^{2}/48}\right).$ (3.30)
-Substitute $z=iy$ and (3.29) into $G(z)$, we obtain
+Substitute $z=iy$ and (3.29) into (3.11), we obtain
 $G(iy)=\frac{1+iy[\frac{1}{6}+\frac{7}{2}C+(\frac{2}{3}-10C)e^{i\psi}]-Cy^{2}(1-e^{i\psi})}{1-(\frac{1}{6}+\frac{13}{2}C)iy-\frac{3}{2}Cy^{2}}.$ (3.31)
@@
 ${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left[\left(\frac{1}{6}+\frac{7}{2}C\right)\mathcal{L}({\bf u}^{n})+\left(\frac{2}{3}-10C\right)\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{6}+\frac{13}{2}C\right)\mathcal{L}({\bf u}^{n+1})\right]$ (3.36)
-$+C\Delta t^{2}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{3}{2}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right],$
-where $C\in[0.018824,0.045589]$, with $C=0.018924$ being the optimal value.
-4 The Implementation of the Implicit TSFO Scheme
-To achieve accelerated convergence, the implicit scheme can be realized through Newton’s iterative method, where the initial guess for each stage uses the result of the two-stage fourth-order explicit scheme [34].
+$+C\Delta t^{2}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{3}{2}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right].$
+The effective range for parameter $C$ is $[0.018824,0.045589]$, within which $C=0.018924$ serves as a favorable choice.
+4 Newton’s Iteration for the Implicit TSFO Scheme
+To achieve accelerated convergence, the implicit TSFO scheme (3.35)-(3.36) can be realized through Newton’s iterative method, where the initial guess for each stage uses the result of the two-stage fourth-order explicit scheme [1].
 Stage 1. The solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$ is computed iteratively as follows:
@@
 $F_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]$ $={\bf u}^{n}+\frac{\Delta t}{4}\left\{\mathcal{L}{({\bf u}^{n})}+\mathcal{L}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]\right\}$
-$\quad-{\frac{\Delta t^{2}}{48}}\left\{\frac{\partial}{\partial t}\mathcal{L}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right\}-({\bf u}^{n+\frac{1}{2}})^{(k)}.$
+$\quad\quad\;\;\;-{\frac{\Delta t^{2}}{48}}\left\{\frac{\partial}{\partial t}\mathcal{L}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right\}-({\bf u}^{n+\frac{1}{2}})^{(k)}.$
 Stage 2. The solution at the next time level ${\bf u}^{n+1}$ is computed iteratively as follows:
@@
 5 Numerical Tests
-This section will validate the convergence order, stability, and competitiveness of the proposed two-stage fourth-order implicit (e.g., Implicit TSFO) scheme (3.35)-(3.36) by applying it to several classic stiff problems, with its fourth-order accuracy being verified through the extrapolation algorithm below,
-$\text{Order}=\frac{\text{ln}[\text{Error}(\Delta t_{1})/\text{Error}(\Delta t_{2})]}{\text{ln}(\Delta t_{1}/\Delta t_{2})},$ (5.1)
-especially, when $\Delta t_{1}=\Delta t$, $\Delta t_{2}=\Delta t/2$, we have
-$\text{Order}_{L^{2}}(\Delta t)=\frac{\text{ln}[\text{Error}_{L^{2}}(\Delta t)/\text{Error}_{L^{2}}(\Delta t/2)]}{\text{ln}2},$ (5.2)
-$\text{Order}_{L^{\infty}}(\Delta t)=\frac{\text{ln}[\text{Error}_{L^{\infty}}(\Delta t)/\text{Error}_{L^{\infty}}(\Delta t/2)]}{\text{ln}2},$ (5.3)
+This section will validate the convergence order, stability, and computational efficiency of the proposed implicit two-stage fourth-order (Implicit TSFO) scheme (3.35)-(3.36) by applying it to several classic stiff problems.
+Its fourth-order accuracy is verified using the following extrapolation algorithm:
+$\text{Order}=\frac{\ln\!\bigl[\text{Error}(\Delta t_{1})/\text{Error}(\Delta t_{2})\bigr]}{\ln\!\bigl(\Delta t_{1}/\Delta t_{2}\bigr)},$ (5.1)
+in particular, when $\Delta t_{1}=\Delta t$ and $\Delta t_{2}=\Delta t/2$, we have
+$\text{Order}_{L^{2}}(\Delta t)=\frac{\ln\!\bigl[\text{Error}_{L^{2}}(\Delta t)/\text{Error}_{L^{2}}(\Delta t/2)\bigr]}{\ln 2},$ (5.2)
+$\text{Order}_{L^{\infty}}(\Delta t)=\frac{\ln\!\bigl[\text{Error}_{L^{\infty}}(\Delta t)/\text{Error}_{L^{\infty}}(\Delta t/2)\bigr]}{\ln 2},$
 where
-$\text{Error}_{L^{2}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{e}\lVert_{L^{2}},\quad\text{Error}_{L^{2}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{c}\lVert_{L^{2}},$ (5.4)
-$\text{Error}_{L^{\infty}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{e}\lVert_{L^{\infty}},\quad\text{Error}_{L^{\infty}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{c}\lVert_{L^{\infty}},$ (5.5)
-${\bf u}_{n}(\Delta t)$ is the numerical solution, ${\bf u}_{e}$ is the exact solution, and ${\bf u}_{c}$ is the reference solution generated by the classical fourth-order explicit Runge-Kutta method with a highly refined time step of $1.0\times 10^{-6}$.
-Additionally, a direct comparison is conducted between the proposed implicit TSFO scheme and the classical fourth-order implicit Runge-Kutta (e.g., Implicit RK4) method, focusing on stable time step size and convergence error, to verify the superiority of the former.
+$\text{Error}_{L^{2}}(\Delta t)$ $=\bigl\|\mathbf{u}_{n}(\Delta t)-\mathbf{u}_{e}\bigr\|_{L^{2}},$ $\text{Error}_{L^{2}}(\Delta t)$ $=\bigl\|\mathbf{u}_{n}(\Delta t)-\mathbf{u}_{c}\bigr\|_{L^{2}},$ (5.3)
+$\text{Error}_{L^{\infty}}(\Delta t)$ $=\bigl\|\mathbf{u}_{n}(\Delta t)-\mathbf{u}_{e}\bigr\|_{L^{\infty}},$ $\text{Error}_{L^{\infty}}(\Delta t)$ $=\bigl\|\mathbf{u}_{n}(\Delta t)-\mathbf{u}_{c}\bigr\|_{L^{\infty}}.$
+Here, $\mathbf{u}_{n}(\Delta t)$ is the numerical solution, $\mathbf{u}_{e}$ is the exact solution, and $\mathbf{u}_{c}$ is a reference solution generated by the classical fourth-order explicit Runge–Kutta method using a highly refined time step of $1.0\times 10^{-6}$.
+Furthermore, to demonstrate the superiority of the proposed implicit TSFO scheme, a direct comparison is conducted with the classical fourth-order implicit Runge–Kutta method (e.g., Implicit RK4), focusing on the allowable stable time step size and the corresponding convergence error.
 5.1 Linear System with Widely Separated Eigenvalues
@@
 Method Implicit TSFO Implicit RK4
-$\Delta t$
-${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
+$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
       $\tau$=1.000000E+00 6.697969115862E-08 –– 8.869206919424E-04 ––
@@
 $\tau/512$=1.953125E-03 9.618514177224E-19 4.0000010755 9.175778879022E-18 4.0000009828
-Table 1: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Linear Systems with Widely Separated Eigenvalues".
+Table 1: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Linear Systems with Widely Separated Eigenvalues
 Method Implicit TSFO Implicit RK4
-$\Delta t$
-${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
+$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
       $\tau$=1.000000E+00 6.697969115862E-08 –– 8.869204367202E-04 ––
@@
 $\tau/512$=1.953125E-03 9.618514177224E-19 4.0000010755 9.175778879022E-18 4.0000009828
-Table 2: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Linear Systems with Widely Separated Eigenvalues".
+Table 2: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Linear Systems with Widely Separated Eigenvalues
 5.2 Robertson’s Chemical Kinetics Problem
@@
 Method Implicit TSFO Implicit RK4
-$\Delta t$
-${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
+$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
         $\tau$=1.000000E-02 9.021072350964E-10 –– 3.854913505270E-09 ––
@@
 $\tau/512$=1.953125E-05 1.851325455787E-20 4.0334732502 1.718639521320E-19 3.9991079765
-Table 3: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Robertson’s Chemical Kinetics Problem".
+Table 3: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Robertson’s Chemical Kinetics Problem
 Method Implicit TSFO Implicit RK4
-$\Delta t$
-${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
+$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
         $\tau$=1.000000E-02 6.379113507692E-10 –– 2.725942952427E-09 ––
@@
 $\tau/512$=1.953125E-05 1.309136414687E-20 4.0334733543 1.215309709126E-19 3.9991079877
-Table 4: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Robertson’s Chemical Kinetics Problem".
+Table 4: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Robertson’s Chemical Kinetics Problem
 5.3 Ozone Decomposition Reaction Problem
@@
 Method Implicit TSFO Implicit RK4
-$\Delta t$
-${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
+$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
        $\tau$=5.0000000E-01 7.304648134165E-05 –– 2.103486775801E-03 ––
@@
 $\tau/512$=9.7656250E-04 1.029202139129E-15 4.0000045767 1.010362911214E-14 4.0000000900
-Table 5: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=1.0$.
+Table 5: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=1.0$
 Method Implicit TSFO Implicit RK4
-$\Delta t$
-${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
+$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
        $\tau$=5.0000000E-01 5.257460149977E-05 –– 1.660009216431E-03 ––
@@
 $\tau/512$=9.7656250E-04 7.717237103705E-16 3.9999985829 7.361066198361E-15 3.9999943997
-Table 6: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=1.0$.
+Table 6: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=1.0$
 Method Implicit TSFO Implicit RK4
-$\Delta t$
-${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
+$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
        $\tau$=5.0000000E-01 1.675891248573E-06 –– 8.943303121509E-06 ––
@@
 $\tau/512$=9.7656250E-04 6.551579640740E-17 3.9997642671 3.459674317438E-16 3.9999898781
-Table 7: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=10.0$.
+Table 7: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=10.0$
 Method Implicit TSFO Implicit RK4
-$\Delta t$
-${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
+$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
        $\tau$=5.0000000E-01 1.627111303401E-06 –– 5.211344689802E-06 ––
@@
 $\tau/512$=9.7656250E-04 6.360313199880E-17 3.9997642539 3.358497322562E-16 3.9999898767
-Table 8: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=10.0$.
+Table 8: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=10.0$
 Method Implicit TSFO Implicit RK4
-$\Delta t$
-${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
+$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
        $\tau$=5.0048554E-01 5.844310485027E-07 –– 3.774174108398E-08 ––
@@
 $\tau/512$=9.7656455E-04 2.211314986480E-17 3.9997632452 1.163453836558E-16 3.9999897963
-Table 9: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=321.8122$.
+Table 9: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=321.8122$
 Method Implicit TSFO Implicit RK4
-$\Delta t$
-${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
+$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
        $\tau$=5.0048554E-01 5.329967588582E-07 –– 3.531284647674E-08 ––
@@
 $\tau/512$=9.7656455E-04 2.016695175240E-17 3.9997632453 1.060908330213E-16 3.9999897949
-Table 10: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=321.8122$.
+Table 10: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=321.8122$
 5.4 Van der Pol Oscillator Problem
@@
 Method Implicit TSFO Implicit RK4
-$\Delta t$
-${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
+$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
         $\tau$=1.000000E-02 4.239641324385E-03 –– 3.224176703077E-03 ––
@@
 $\tau/512$=1.953125E-05 6.334338803941E-15 4.0005276704 6.044905950897E-14 4.0000477364
-Table 11: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Van der Pol Oscillator Problem".
+Table 11: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Van der Pol Oscillator Problem
 Method Implicit TSFO Implicit RK4
-$\Delta t$
-${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
+$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
         $\tau$=1.000000E-02 4.239531568484E-03 –– 3.224093052535E-03 ––
@@
 $\tau/512$=1.953125E-05 6.334173311281E-15 4.0005276704 6.044748020038E-14 4.0000477364
-Table 12: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Van der Pol Oscillator Problem".
-Numerical results confirm that the proposed two-step fourth-order implicit (Implicit TSFO) scheme effectively solves stiff problems using large time steps without compromising fourth-order accuracy.
-When compared to the classical implicit Runge-Kutta (Implicit RK4) method, our scheme reduces the numerical error by about an order of magnitude under identical time steps and offers superior stability that enables even larger step sizes.
+Table 12: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Van der Pol Oscillator Problem
+The numerical experiments confirm that the proposed implicit TSFO time discretization scheme successfully achieves fourth-order temporal accuracy in only two computational stages.
+The results also demonstrate the robust stability of the proposed algorithm when solving stiff problems with relatively large time steps, thereby validating its effectiveness for stiff systems.
+In comparison to the classical fourth-order implicit Runge-Kutta method, the present scheme produces smaller errors under identical time step sizes, typically reducing the error by an order of magnitude (reference solutions are computed using the explicit fourth-order Runge-Kutta method).
+Moreover, it maintains stability under even larger time steps, highlighting its superior numerical stability.
+The source code for all test cases presented is publicly available in the accompanying GitHub repository at: https://github.com/ZhixinHuo/Implicit-TSFO.
 6 Conclusions and future work
-This paper presents a novel two-stage fourth-order implicit scheme.
-We rigorously establish its fourth-order accuracy through mathematical analysis and systematically examine the scheme’s stability, deriving the conditions under which A-stability can be achieved.
-A detailed implementation procedure based on Newton iteration is provided.
-The effectiveness of the proposed method is validated through numerical experiments on several classical strongly stiff benchmark cases.
-Comparisons with classical explicit and implicit fourth-order Runge-Kutta methods demonstrate that the proposed scheme possesses superior stability regions and smaller numerical errors.
-In future work, we will combine this implicit scheme with its explicit counterpart—a spatiotemporally coupled two-stage fourth-order method—to solve fluid dynamics problems characterized by strongly stiff source terms, such as those encountered in chemical reactions.
-Compared to other types of implicit schemes, the proposed method incorporates temporal derivatives of physical quantities, which is expected to enhance both computational accuracy and compactness of the scheme.
+This study fully leverages the time derivatives of physical quantities to construct a implicit TSFO temporal discretization scheme through rigorous mathematical derivation.
+A systematic theoretical analysis provides sufficient conditions for the scheme to achieve A-stability, and a corresponding Newton iteration process is proposed to accelerate convergence.
+The proposed algorithm is validated on a series of classical stiff benchmark problems, demonstrating its robustness and competitiveness.
+Numerical experiments show that the scheme achieves fourth-order temporal accuracy using only two stages.
+Compared with classical fourth-order implicit Runge-Kutta methods, it not only allows larger stable time steps but also yields smaller convergence errors.
+This study establishes a solid theoretical foundation and a concrete implementation pathway for extending TSFO time discretization methods from existing explicit frameworks [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] to implicit frameworks.
+Its significance and necessity lie not only in extending the applicability of TSFO methods from idealized homogeneous hydrodynamic problems to realistic ones with stiff source terms, but also in unlocking the inherent potential of Lax-Wendroff-based GRP solvers to naturally handle stiff source terms within the TSFO framework.
+The ability of TSFO methods to achieve fourth-order temporal accuracy with only two stages stems from their utilization of the time derivatives of physical quantities, which are precisely provided by Lax-Wendroff-based GRP solvers.
+Compared to traditional Riemann solvers, these advanced solvers can compute temporal derivatives that simultaneously integrate the physical effects of both fluxes and source terms within a unified framework through the governing equations, thereby effectively avoiding the numerical errors introduced by operator-splitting approaches.
+Future work will integrate the existing explicit TSFO time discretization method [1] with the implicit TSFO scheme proposed in this paper, and further combine it with a more compact HWENO spatial discretization method based on GRP solver [5].
+This integrated framework aims to solve hydrodynamic problems involving stiff source terms, targeting the simulation of complex, multi-scale, and multi-physics phenomena in real-world systems, such as chemically reacting flows, radiation hydrodynamics, combustion dynamics, and other stiff physical systems.
 7 Acknowledgments
@@
 26A110007), the Young Talents Fund of Henan Province (Grant No.
-252300423500), the "Double First-Class" Project of the School of Geomatics, Henan Polytechnic University (Grant No.
+252300423500), the Double First-Class Project of the School of Geomatics of Henan Polytechnic University (Grant No.
 BSJJ202306), and the Doctoral Startup Foundation of Henan Polytechnic University (Grant No.